### v1

# Graphon branching processes and fractional isomorphism

Abstract
In their study of the giant component in inhomogeneous random graphs, Bollobás, Janson, and Riordan introduced a class of branching processes parametrized by a possibly unbounded graphon. We prove that two such branching processes have the same distribution if and only if the corresponding graphons are fractionally isomorphic, a notion introduced by Grebík and Rocha.
A different class of branching processes was introduced by Hladký, Nachmias, and Tran in relation to uniform spanning trees in finite graphs approximating a given connected graphon. We prove that two such branching processes have the same distribution if and only if the corresponding graphons are fractionally isomorphic up to scalar multiple. Combined with a recent result of Archer and Shalev, this implies that if uniform spanning trees of two dense graphs have a similar local structure, they have a similar scaling limit.
As a side result we give a characterization of fractional isomorphism for graphs as well as graphons in terms of their connected components.
1 Introduction
1.1 Bollobás–Janson–Riordan branching processes
Branching processes related to graphs appear most prominently in the study of the giant component in the Erdős–Rényi random graph $\mathbb{G}(n,\frac{d}{n})$. While the original approach, starting with the seminal work of Erdős and Rényi [10], used enumerative techniques, in the 1990s it was realized that the local structure of $\mathbb{G}(n,\frac{d}{n})$ can be approximated by the Galton–Watson branching process $\mathfrak{X}_{d}$ with offspring distribution $\mathsf{Poi}(d)$. The size of the giant component can be expressed in terms of $\mathfrak{X}_{d}$, too. As was kindly pointed out to us by Tomasz Łuczak, this idea first appeared in [18]. To make the statement more precise, we have the following: for a branching process $\mathfrak{X}$ and $k\in\mathbb{N}_{0}$, we write $(\mathfrak{X})_{\restriction k}$ for the $k$-ball of $\mathfrak{X}$ around the root.
Fact 1.1.
Let $d\geq 0$ be given. If $k\in\mathbb{N}$ and $T$ is a rooted graph whose vertices are all at distance at most $k$ from the root then the probability that the $k$-ball around vertex 1 in $\mathbb{G}(n,\frac{d}{n})$ is isomorphic to $T$ converges to $\mathbf{P}[(\mathfrak{X}_{d})_{\restriction k}\cong T]$, as $n\to\infty$. The order of the largest connected component in $\mathbb{G}(n,\frac{d}{n})$ is $(s+o_{\mathrm{p}}(1))n$, where $s$ is the survival probability of $\mathfrak{X}_{d}$. (The term $o_{\mathrm{p}}(1)$ converges to 0 in probability, as $n\to\infty$.)
A significant extension of the Erdős–Rényi random graph was introduced by Bollobás, Janson, and Riordan in [7], with further important contributions to the subject in [5, 6, 17]. For our purposes, it suffices to introduce a slightly less general version of their model. Let $(X,\mathcal{B})$ be a standard Borel space endowed with a Borel probability measure $\mu$. A kernel is a symmetric measurable function $W:X^{2}\rightarrow[0,\infty)$. The degree of $x\in X$ in $W$ is given by $\deg_{W}(x)=\int_{y\in X}W(x,y)\mathsf{d}\mu(y)$. The maximum degree and minimum degree of $W$ are defined as the essential supremum and the essential infimum of the degree, $\operatorname{deg^{min}}(W)=\operatorname{ess\,inf}_{x}\deg_{W}(x)$, $\operatorname{deg^{max}}(W)=\operatorname{ess\,sup}_{x}\deg_{W}(x)$. When $\|W\|_{1}<\infty$, $\operatorname{deg^{max}}(W)<\infty$, $\|W\|_{\infty}<\infty$, or $\|W\|_{\infty}\leq 1$, we call $W$ an $L^{1}$-kernel, bounded-degree kernel, $L^{\infty}$-kernel, or graphon, respectively. The sparse inhomogeneous random graph $\mathbb{G}(n,\frac{W}{n})$ is defined on the vertex set $[n]$ as follows. First, sample elements $x_{1},\ldots,x_{n}\in X$ independently according to $\mu$. For each $\{i,j\}\in\binom{[n]}{2}$ independently, include $ij$ as an edge of $\mathbb{G}(n,\frac{W}{n})$ with probability $\min\{1,\frac{W(x_{i},x_{j})}{n}\}$. In particular, if $W\equiv d$ for some $d\in[0,\infty)$, we get the Erdős–Rényi random graph $\mathbb{G}(n,\frac{d}{n})$. To introduce a counterpart of Fact 1.1 for $\mathbb{G}(n,\frac{W}{n})$, we define the inhomogeneous Galton–Watson branching process $\mathfrak{X}_{W}$ as a multitype Galton–Watson branching process with type space $X$. The type of the root is selected according to the distribution $\mu$. Each particle of type $x\in X$ has children whose number and types follow a Poisson point process on $X$ with intensity $W(x,y)\mathsf{d}\mu(y)$. Observe that when $W$ is an $L^{1}$-kernel, $(\mathfrak{X}_{W})_{\restriction k}$ is finite almost surely for every $k\in\mathbb{N}$. The counterpart of Fact 1.1 for $\mathbb{G}(n,\frac{W}{n})$, proved in [7], then reads as follows.
Fact 1.2.
Let $W:X^{2}\rightarrow[0,\infty)$ be an $L^{1}$-kernel. If $k\in\mathbb{N}$ and $T$ is a rooted graph whose vertices are all at distance at most $k$ from the root, the probability that the $k$-ball around vertex 1 in $\mathbb{G}(n,\frac{W}{n})$ is isomorphic to $T$ converges to $\mathbf{P}[(\mathfrak{X}_{W})_{\restriction k}\cong T]$, as $n\to\infty$. The order of the largest component in $\mathbb{G}(n,\frac{W}{n})$ is $(\gamma(W)+o_{\mathrm{p}}(1))n$, where $\gamma(W)$ is the survival probability of $\mathfrak{X}_{W}$.
The survival probability $s$ of $\mathfrak{X}_{d}$ is known to be 0 for $d\leq 1$, and to be given as the unique solution $s\in(0,1)$ of the equation $1-s=\exp(-ds)$ for $d>1$ (see e.g. [2, §10.4]). In [7], an inhomogeneous counterpart is obtained for $\mathfrak{X}_{W}$. However, the survival probability $\gamma(W)$ is concealed in a solution of a function-valued generalization of the above real-valued equation. Even for rather simple $L^{1}$-kernels $W$, this functional equation is not tractable. There are other quantities based on Bollobás–Janson–Riordan branching processes which are even more mysterious. For example, (a generalization of) the following setting is considered in [14]. Suppose $(H_{n})_{n}$ is a sequence of connected graphs converging (in the sense of dense graph convergence) to a connected [^1] A kernel $W$ is connected if $\int_{A\times(X\setminus A)}W>0$ for every $A\subset X$ with $\mu(A)\in(0,1)$. See Section 2.1 for details. kernel $W$. On each edge of each $H_{n}$, put an independent weight chosen uniformly from $[0,1]$, and consider the minimum spanning tree $T_{n}$ of $H_{n}$. Then the total weights of $T_{n}$ converge in probability to a constant $\kappa(W)=\int_{t=0}^{\infty}\sum_{k=1}^{\infty}\frac{\mathbf{P}[|\mathfrak{X}_{t\cdot W}|=k]}{k}$. Expressing $\kappa(W)$ will be extremely challenging if not impossible. Note that the simplest case $W\equiv d\in[0,\infty)$ can be computed explicitly, see [14, §6.3].
Our main result concerning Bollobás–Janson–Riordan branching processes characterizes bounded-degree kernels $U$ and $W$ for which $\mathfrak{X}_{U}$ and $\mathfrak{X}_{W}$ have the same distribution. We emphasize that by saying that two branching processes have the same distribution, we mean they produce the same distribution on isomorphism classes of rooted (and possibly infinite) trees. [^2] Recall that two distribution $\mathfrak{T}$ and $\mathfrak{D}$ on rooted infinite trees are equal if for every $k\in\mathbb{N}$, $\mathfrak{T}_{\restriction k}$ and $\mathfrak{D}_{\restriction k}$ have the same distribution. In particular, we disregard the labeling of the particles by their types.
Fractional isomorphism for graphs is a concept introduced in 1986 in [25] as a relaxation of graph isomorphism. Among several equivalent definitions, we give a rather intuitive one stated in terms of the color refinement algorithm, which we describe below. This definition will naturally generalize to the definition of fractional isomorphism of kernels that we will primarily work with. In Section 5 we will see a different but equivalent definition of fractional isomorphism for graphs. The color refinement algorithm iteratively colors the vertices of a finite graph. It starts by giving each vertex the same color. Then, in each step it recolors vertices of the same color in the following way: if their respective neighbors have the same colors (with multiplicities), they will have the same color afterwards, otherwise, they get distinct colors. When vertices are recolored, they only get colors which were not already used to color different vertices. The algorithm stops as soon as no vertices need to be recolored by the above rules. Two finite graphs of the same order are fractionally isomorphic if the color refinement algorithm applied to their disjoint union yields a vertex coloring where each color class has the same number of vertices in each of the two graphs. In particular, any two $d$-regular $n$-vertex graphs are fractionally isomorphic, as the algorithm terminates immediately with a uniform coloring.
Remarkably, the above definition of fractional isomorphism is equivalent by results from [8, 9, 23] to the existence of a doubly stochastic matrix $P$ so that for the adjacency matrices $A_{H}$ and $A_{G}$ of $H$ and $G$ we have $A_{H}=P^{-1}A_{G}P$, and also to having equal counts of each tree in $H$ and in $G$.
The above notions of fractional isomorphism were translated to graphons by Grebík and Rocha in [11], where they also showed that all these notions are equivalent in the graphon setting. In Section 2.2, we recall their definition of fractional isomorphism for graphons which corresponds to the notion for graphs given in terms of iterated degree sequences. In fact, we recall their definition in a slightly more general setting where we work with a generalized version of graphons that does not require boundedness and permits asymmetry; we call these objects akernels. In particular, we call a measurable function $W:X^{2}\rightarrow[0,\infty)$ a bounded-degree akernel if $\operatorname{deg^{max}}(W)<\infty$ and an $L^{\infty}$-akernel if $\|W\|_{\infty}<\infty$. The definition of the branching process $\mathfrak{X}_{W}$ is sensible for bounded-degree akernels as well. That is, the assumption of symmetry is not needed. In particular, it is the slices $W(x,\cdot)$ rather than $W(\cdot,x)$ that appear both in the definition of bounded degree akernels as well as in the definition of offspring generation in $\mathfrak{X}_{W}$. The generalization to the asymmetric setting may not seem important in the context of the original setting of Bollobás, Janson, and Riordan. However, this asymmetric version will play a key role in the proof of Theorem 1.7.
Theorem 1.3.
Suppose that $U$ and $W$ are bounded-degree akernels. Then $\mathfrak{X}_{U}$ and $\mathfrak{X}_{W}$ have the same distribution if and only if $U$ and $W$ are fractionally isomorphic.
As we show in Proposition 3.1, the ()($\Leftarrow$)( ⇐ ) direction of Theorem 1.3 is easy. The ()($\Rightarrow$)( ⇒ ) direction, which is the main challenge of the theorem, is proved in Section 3.
In particular, Theorem 1.3 allows us to say that the survival probabilities $\gamma(W_{1})$ and $\gamma(W_{2})$ are equal whenever $W_{1}$ and $W_{2}$ are fractionally isomorphic, even though we will be typically unable to determine their values. Since fractional isomorphism of $W_{1}$ and $W_{2}$ easily implies fractional isomorphism of $t\cdot W_{1}$ and $t\cdot W_{2}$ for every $t\geq 0$, we also get $\kappa(W_{1})=\kappa(W_{2})$. This is in particular useful for $d$-regular kernels. Recall that a kernel $W$ is $d$-regular (for $d\in[0,\infty)$) if for $\mu$-almost every $x\in X$ we have $\deg_{W}(x)=d$. The class of all $d$-regular kernels may seem quite complicated, but the above implies that any parameter derived from $\mathfrak{X}_{W}$ is constant on $d$-regular kernels, and in particular is equal to its value for $\mathfrak{X}_{d}$, where it may be tractable computationally, as was the case with $\gamma(\mathfrak{X}_{d})$ and $\kappa(\mathfrak{X}_{d})$.
1.2 Local structure of uniform spanning trees
Our second main result is related to branching processes arising from the study of the uniform spanning tree. Recall that the uniform spanning tree of a connected graph $G$ is the uniform measure on spanning trees of $G$. While much of the study of uniform spanning trees is concerned with large but sparse graphs with lattice structure, here we are concerned with dense graphs. The first result that relates to the uniform spanning tree of dense graphs is by Kolchin [19] and Grimmett [12]. These results are best expressed using the notion of local limit (also known as the Benjamini–Schramm limit), which we now recall.
Definition 1.4.
Suppose that $(F_{n})_{n}$ is a sequence of graphs. Let $\mathfrak{G}$ be a probability distribution on isomorphism classes of rooted (possibly infinite) graphs with finite degrees. We say that $(F_{n})_{n}$ converges locally to $\mathfrak{G}$ if for every radius $r\in\mathbb{N}$, the distribution of the $r$-neighborhood of a uniformly chosen random vertex in $F_{n}$ converges to the distribution of the $r$-neighborhood around the root in $\mathfrak{G}$.
The result of Kolchin and Grimmett says that for each $r\in\mathbb{N}$ the uniform spanning trees of $K_{n}$ converge, as $n\to\infty$, locally to the Galton–Watson branching process with offspring distribution $\mathsf{Poi}(1)$, conditioned on survival. It is well-known that such a branching process can be constructed by taking a one-way infinite path rooted at its endvertex and attaching a $\mathsf{Poi}(1)$ Galton–Watson branching process to each vertex of the aforementioned one-way infinite path. [^3] Let us include details. Section 5 of [16] contains a construction of a modified Galton–Watson branching process by introducing a size-biased version of the offspring distribution. Since we started with a $\mathsf{Poi}(1)$ Galton–Watson branching process, the modified Galton–Watson branching process has offspring distribution in which each particle produces $k$ offspring with probability $k\mathbf{P}[\mathsf{Poi}(1)=k]$. It is easy to see that the size-biased $\mathsf{Poi}(1)$ distribution is equal to $1+\mathsf{Poi}(1)$, and hence corresponds to the above construction with a one-way infinite path. Theorem 22.2 in [16] shows that the modified Galton–Watson branching process is indeed the $\mathsf{Poi}(1)$ Galton–Watson branching process conditioned on survival. This corresponds to the branching process $\mathfrak{U}_{W}$ defined below, when we take $W\equiv 1$.
The above result was generalized from complete graphs to sequences converging to a graphon by Hladký, Nachmias, and Tran [13]. Let us introduce their definition in a more general setting of $L^{1}$-kernels. We say that an $L^{1}$-kernel $W$ is nondegenerate if $\deg_{W}(x)>0$ for almost every $x\in X$.
Definition 1.5.
Given a nondegenerate $L^{1}$-kernel $W$, we define a multitype Galton–Watson branching process $\mathfrak{U}_{W}$ with type space $\{(\mathsf{anc},x),(\mathsf{oth},x)\;:\>x\in X\}$ as follows. Here, “$\mathsf{anc}$” stands for “ancestral” and “$\mathsf{oth}$” stands for “other”.
  1. (i)
    The root has type $(\mathsf{anc},x)$, where the distribution of $x$ is $\mu$.
  2. (ii)
    If a particle has type $(\mathsf{oth},x)$, then its children are $\{(\mathsf{oth},x_{1}),\ldots,(\mathsf{oth},x_{k})\}$ where $\{x_{1},\ldots,x_{k}\}$ are the Poisson point process on $X$ with intensity $\frac{W(x,y)}{\deg_{W}(y)}$ at $y\in X$. In particular, the number of children of this particle has distribution $\mathsf{Poi}\left(\int_{y}\frac{W(x,y)}{\deg_{W}(y)}\right)$.
  3. (iii)
    If a particle has type $(\mathsf{anc},x)$, then its children are $\{(\mathsf{anc},x_{0}),(\mathsf{oth},x_{1}),\ldots,(\mathsf{oth},x_{k}))\}$ where $\{x_{1},\ldots,x_{k}\}$ are the Poisson point process on $X$ with the same intensity as above and $x_{0}$ is an independent element which is distributed according to the probability measure on $X$ that has density at each $y\in X$ equal to $\frac{W(x,y)}{\deg_{W}(x)}$. In particular, the number of children of this particle has distribution $1+\mathsf{Poi}\left(\int_{y}\frac{W(x,y)}{\deg_{W}(y)}\right)$.
Figure 1: The branching process $\mathfrak{U}_{W}$ from Definition 1.5. Let $\mathfrak{U}^{-}_{W}$ be obtained from $\mathfrak{U}_{W}$ by removing the unique $\mathsf{anc}$-vertex neighboring the root and the whole subtree appended to it. The complement of $\mathfrak{U}_{W}^{-}$ is called the ancestral part of $\mathfrak{U}_{W}$. The one-way infinite path rooted at the root of the tree consisting of $\mathsf{anc}$-vertices is called the ancestral path.
In Figure 1 we give an example of the branching process $\mathfrak{U}_{W}$. In that figure we also introduce a branching process $\mathfrak{U}_{W}^{-}$ and the ancestral part of $\mathfrak{U}_{W}$.
The main result of [13] then reads as follows.
Theorem 1.6.
Let $W:X^{2}\rightarrow[0,\infty)$ be an arbitrary nondegenerate graphon. Suppose that $(G_{n})_{n}$ is a sequence of connected graphs of growing orders that converges to $W$ in the cut distance. The uniform spanning trees of $G_{n}$ converge locally to $\mathfrak{U}_{W}$, as $n\to\infty$. In particular, if $r\in\mathbb{N}$ and $T$ is an arbitrary rooted graph whose all vertices are at distance at most $r$ from the root then we have that the probability that the $r$-ball around a randomly selected vertex the uniform spanning tree of $G_{n}$ is isomorphic to $T$ converges to $\mathbb{P}[(\mathfrak{U}_{W})_{\restriction r}\cong T]$, as $n\to\infty$.
Like in Theorem 1.3, it can be checked that $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have the same distribution if $U$ and $W$ are fractionally isomorphic. There is however another operation which is easily seen to preserve the distribution of $\mathfrak{U}_{W}$, namely multiplying by a scalar. Our main result says that there are no further ways of creating a connected kernel $Z$ with the same distribution as $\mathfrak{U}_{Z}$. Hence, we say that two kernels $U$ and $W$ are projectively fractionally isomorphic if there exists a constant $t>0$ so that $U$ and $tW$ are fractionally isomorphic. This is equivalent to $\frac{U}{\|U\|_{1}}$ and $\frac{W}{\|W\|_{1}}$ being fractionally isomorphic. We state the theorem only for connected kernels. A full characterization including disconnected kernels is given as Theorem 6.1 in Section 6.
Theorem 1.7.
Suppose that $U$ and $W$ are connected $L^{\infty}$-kernels with $\operatorname{deg^{min}}(U),\operatorname{deg^{min}}(W)>0$. Then $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have the same distribution if and only if $U$ and $W$ are projectively fractionally isomorphic.
Theorem 1.7 has an interesting corollary which connects the local limit and the scaling limit of uniform spanning trees on dense graphs. We briefly recall the latter concept. Aldous [1] famously constructed a certain random metric space, which he called ‘the continuum random tree’ $\mathcal{T}$. He showed that $\mathcal{T}$ is the Gromov–Hausdorff–Prohorov limit of uniform spanning tree on $K_{n}$ when each edge is given length $n^{-1/2}$, as $n\to\infty$. This type of convergence is called the ‘scaling limit’. Since then, $\mathcal{T}$ has been shown to be the scaling limit of many other sequences of uniform spanning trees. For dense graphs, a recent result of Archer and Shalev [3] asserts that in the setting of Theorem 1.6 (with an additional assumption that $W$ is connected), the scaling limit of the uniform spanning trees is $c_{W}\cdot\mathcal{T}$, where the rescaling constant $c_{W}$ is defined by (see [3, Equation (1)])
$c_{W}:=\sqrt{\frac{\int_{x}\deg_{W}(x)^{2}\mathsf{d}\mu(x)}{\|W\|_{1}^{2}}}\;.$ (1)
Crucially for us, if $U$ and $W$ are projectively fractionally isomorphic, then $c_{U}=c_{W}$. We give a proof in Lemma 2.4.
The aforementioned connection between the local and the scaling limit applies to a class of graphs which do not have sparse cuts. More precisely, for $\rho>0$ we say that a graph $G$ is $\rho$-robust if for every $U\subset V(G)$ we have $e_{G}(U,V(G)\setminus U)\geq\rho|U||V(G)\setminus U|$.
Corollary 1.8.
Let $\rho>0$ be an arbitrary constant. Suppose that $(G_{n})_{n}$ is a sequence of $\rho$-robust graphs of growing orders $v(G_{n})$. Suppose that the distribution of the uniform spanning trees of $G_{n}$ converges locally, as $n\to\infty$. Then the uniform spanning tree on $G_{n}$ whose each edge is given length $v(G_{n})^{-1/2}$ converges in the sense of the scaling limit, as $n\to\infty$.
To prove Corollary 1.8, let $\mathcal{U}$ be all those graphons that arise as accumulation points of the sequence $(G_{n})_{n}$ with respect to the cut distance topology. Theorems 1.6 and 1.7 tell us that $\mathcal{U}$ is contained within one equivalence class of projective fractional isomorphism. In particular, the constant defined in (1) is universal for $\mathcal{U}$, say $c_{\mathcal{U}}$. Further, the assumption of robustness implies that all the graphons in $\mathcal{U}$ are connected, which is necessary for the result of Archer and Shalev to apply. Thus, the result of Archer and Shalev tells us that the uniform spanning tree on $G_{n}$ with the edge lengths rescaled by $v(G_{n})^{-1/2}$ converges in the sense of the scaling limit to $c_{\mathcal{U}}\cdot\mathcal{T}$, as $n\to\infty$.
1.3 A factorization result for fractional isomorphism
In Section 5 we state and prove Theorem 5.1, which is a factorization result for fractional isomorphism of disconnected graphons. Roughly, it says that two graphons, $U$ and $W$, are fractionally isomorphic if and only if, for each connected graphon $\Gamma$, the total measure of connected components which are fractionally isomorphic to $\Gamma$ is the same in both $U$ and $W$. This result is an addition to the theory worked out in [11] and does not directly concern graphon branching processes. The reason we include it here is that it allows us to extend Theorem 1.7 to disconnected graphons as we do in Theorem 6.1. We also include Theorem 5.2, which is a graph counterpart of Theorem 5.1. While Theorem 5.2 is not needed for our results about branching processes, we consider it an important contribution to the theory of graph fractional isomorphism.
1.4 Organization of the paper
Section 2 contains preliminaries for the proof. In Sections 2.12.3 we provide tools that will be useful for our proof of Theorem 1.3: general background on graphons, fractional isomorphism of graphons (mostly from [11], with minor adjustments), and a suitable version of the Stone–Weierstraß Theorem. New material appears in Section 3.1. Specifically, Section 3.1 introduces further useful branching processes and establishes identities which are useful for the proof of Theorem 1.3. The now short proof of Theorem 1.3 is given in Section 3.
It turns out that the proofs of our results about branching processes $\mathfrak{U}_{W}$, Theorem 1.7 and its generalization for disconnected kernels given in Theorem 6.1, require more preparations. In Section 4.2 we introduce a transformation $W^{\dagger}$ of a graphon $W$ which is suitable to study part $\mathfrak{U}_{W}^{-}$ of the branching process $\mathfrak{U}_{W}$. In particular, the main result of Section 4.2.1 says that $\mathfrak{U}_{W}^{-}$ goes extinct almost surely. We will also need to recall the theory of discrete time Markov chains on uncountable measure spaces, we do that in Section 4.1. In Section 5 we then state the factorization theorem for fractional isomorphism needed for extending Theorem 1.7 to Theorem 6.1. With these preparations, Section 6 contains the proof of Theorem 6.1.
1.5 Acknowledgments
We thank Ellie Archer and Jan Grebík for discussions.
2 Notation, preliminaries, and tools for proof of Theorem 1.3
We write =$\mathbb{N}=\{1,2,\ldots\}$, $\mathbb{N}_{0}=\{0,1,2,\ldots\}$, =$\mathbb{N}_{\infty}=\{1,2,\ldots,\infty\}$, and $\mathbb{N}_{0,\infty}=\{0,1,2,\ldots,\infty\}$. For $n\in\mathbb{N}_{0}$ we write $[n]=\{1,2,\ldots,n\}$ and $[n]_{0}=\{0,1,2,\ldots,n\}$. We write $\mathbf{P}$ and $\mathbf{E}$ for probability and expectation, respectively. Given a measure space $(X,\mathcal{B},\mu)$ and $p\in[1,\infty]$, we write $L^{p}(X,\mathcal{B},\mu)$ for the space of all $\mathcal{B}$-measurable functions of finite $L^{p}$-norm with respect to $\mu$. When $\mathcal{B}$ or $\mu$ are clear from the context, we write just $L^{p}(X,\mu)$ or $L^{p}(X)$.
After introducing connected components of kernels in Section 2.1, in Section 2.2 we introduce the notion of fractional isomorphism which is used in the statement of our main results. Objects such as $\mathbb{P}^{c}$ and $\nu_{W}$ introduced in this section play a key role in the main proofs. The proof of Theorem 1.3 relies on a version of the Stone–Weierstraß Theorem as we introduce it in Section 2.3.
Notation we introduce in the paper is also summarized in Appendix A (available in the arXiv version).
2.1 Graph limits basics
We now recall the notion of connected kernels (sometimes also called irreducible kernels) and of connected components of kernels. This notion was thoroughly studied in [15] and has since become standard.
Definition 2.1.
Let $I$ be a finite or countable set which does not contain 00. For a kernel $W$ we say that a decomposition $X=\Lambda_{0}\sqcup\bigsqcup_{i\in I}\Lambda_{i}$ into finitely or countably many sets is a decomposition into connected components of $W$ with isolated elements $\Lambda_{0}$ if we have the following
  1. (i)
    For $\mu$-almost every $x\in\Lambda_{0}$ we have $\deg_{W}(x)=0$.
  2. (ii)
    For every $i\in I$ we have $\mu(\Lambda_{i})>0$. Further, for every $B\subset\Lambda_{i}$ with $\mu(B)\in(0,\mu(\Lambda_{i}))$ we have $\int_{B\times(\Lambda_{i}\setminus B)}W>0$.
  3. (iii)
    For every $i\in I$ we have $\int_{\Lambda_{i}\times(X\setminus\Lambda_{i})}W=0$.
The sets $\Lambda_{i}$ ($i\in I$) are called the connected components of $W$.
It is shown in [15] that a decomposition into connected components exists and is unique modulo nullsets. Further, if $W$ is nondegenerate, then $\Lambda_{0}$ is obviously a nullset. Also, we say that $W$ is connected if $\Lambda_{0}$ is a nullset and $|I|=1$, a definition consistent with the one we included in Footnote 1.
2.2 Fractional isomorphism
In this section, we give a definition of fractional isomorphism for bounded-degree akernels. The theory developed by Grebík and Rocha [11] gives five equivalent definitions of fractional isomorphism for graphons, and it straightforwardly generalizes in its entirety from graphons to kernels of bounded $L^{\infty}$-norm. We use one of those five definitions and generalize it to the slightly more general setting of bounded-degree akernels.444 An example of a bounded-degree kernel $W:(0,1)\rightarrow[0,\infty)$ which does not have finite $L^{\infty}$-norm is as follows. For $x,y$ such that $x,y\in(2^{-k},2^{-(k-1)})$ for some $k\in\mathbb{N}$, define $W(x,y)=2^{k}$. Otherwise, set $W(x,y)=0$. However, we do not claim that generalizations of their other definitions are equivalent or even well-defined. Later, in Section 4.4 we recall a different characterization of fractional isomorphism for the purposes of the proof of Theorem 1.7.
To prepare for our definition of fractional isomorphism for bounded-degree akernels, we will do the following. First, we introduce a compact metric space $\mathbb{P}^{c}$ whose elements represent encapsulations of iterated degree information. Second, for each bounded-degree akernel $W$ we define a map $i_{W}\colon X\to\mathbb{P}^{c}$, where $i_{W}(x)$ captures the iterated degree information of $x$ in $W$. Finally, we assign a measure $\nu_{W}$ on $\mathbb{P}^{c}$ to each $W$; this measure plays a key role in our definition of fractional isomorphism.
We recall a general definition of a pushforward measure. Given a measurable map $f:A\to B$ between measurable spaces and a measure $\alpha$ on $A$, the pushforward of $\alpha$ via $f$, denoted by $(f)_{*}\alpha$, is a measure on $B$ satisfying $(f)_{*}\alpha(S):=\alpha(f^{-1}(S))$ for every measurable $S\subset B$.
Let $c\in\mathbb{R}_{+}$. For a compact metric space $K$ we write $\mathcal{M}_{\leq c}(K)$ for the set of all Borel measures on $K$ with total mass at most $c$ and we define $\mathcal{M}_{=c}(K)$ analogously. We equip both $\mathcal{M}_{\leq c}(K)$ and $\mathcal{M}_{=c}(K)$ with the topology of weak convergence. We recall the basics and refer to Sections 1 and 2 of [4] for details. The weak topology of measures is metrizable and thus also characterized by convergent sequences; a sequence $\zeta_{1},\zeta_{2},\ldots\in\mathcal{M}_{\leq c}(K)$ converges weakly to a measure $\zeta$ if for every bounded continuous function $f:K\rightarrow\mathbb{R}$ we have
$\lim_{n}\int f\mathsf{d}\zeta_{n}=\int f\mathsf{d}\zeta\;.$ (2)
Recall also Prokhorov’s Theorem, which states that $\mathcal{M}_{\leq c}(K)$ is compact.
To construct $\mathbb{P}^{c}=\mathbb{P}^{c}_{\infty}$ we shall recursively define collections $\{L_{k}^{c}\}_{k\in\mathbb{N}_{0}}$ and $\{\mathbb{P}_{k}^{c}\}_{k\in\mathbb{N}_{0}}$ of spaces as well as canonical projections $p_{\ell,k}^{c}$ from $\mathbb{P}_{k}^{c}$ to $\mathbb{P}_{\ell}^{c}$ for $k>\ell$.
Let $L_{0}^{c}=\{\star\}$ be the one-point space. For $k\in\mathbb{N}_{0}$, let
$\mathbb{P}_{k}^{c}:=\left\{\left.\alpha\in\prod_{i=0}^{k}L_{i}^{c}\ \right|\ \forall\ell\in[k-1]:\ \alpha(\ell)=(p^{c}_{\ell-1,\ell})_{*}\alpha(\ell+1)\right\},$ (3)
where the condition $\forall\ell\in[k-1]$ is vacuous for $k=0,1$. Let
$p_{\ell,k}^{c}\colon\mathbb{P}_{k}^{c}\to\mathbb{P}_{\ell}^{c}$ (4)
be the canonical projection for $\ell\in[k]_{0}$, and let
$L_{k+1}^{c}:=\mathcal{M}_{\leq c}(\mathbb{P}_{k}^{c})\;.$ (5)
We note that the canonical projections $p_{\ell,k}^{c}$ are well defined and we obtain inductively that the spaces $\mathbb{P}_{k}^{c}$ and $L_{k+1}^{c}$ are compact.
The canonical projections, for $n,k\in\mathbb{N}_{0}$ with $n\leq k$, naturally give rise to the pushforwards $(p_{\ell,k}^{c})_{*}\colon L_{k+1}^{c}\to L_{\ell+1}^{c}$. Finally, let
$\mathbb{P}^{c}=\mathbb{P}^{c}_{\infty}=\left\{\left.\alpha\in\prod_{i=0}^{\infty}L_{i}^{c}\right|\ \forall k\in\mathbb{N}:\ \alpha(k)=(p^{c}_{k-1,k})_{*}\alpha(k+1)\right\}$ (6)
and for $k\in\mathbb{N}_{0}$ let
$p_{k,\infty}^{c}\colon\mathbb{P}_{\infty}^{c}\to\mathbb{P}_{k}^{c}\;$ (7)
be the corresponding canonical projections. The space $\mathbb{P}^{c}$ is compact by Tychonoff’s Theorem. We write $\mathcal{B}(\mathbb{P}^{c})$ and $\mathcal{B}(\mathbb{P}_{n}^{c})$ for the set of of Borel sets on $\mathbb{P}^{c}$ and $\mathbb{P}_{n}^{c}$, respectively.
For each $k\in\mathbb{N}_{\infty}$ and $\alpha\in\mathbb{P}_{k}^{c}$, we have $\alpha(1)(\mathbb{P}^{c}_{0})=\alpha(2)(\mathbb{P}^{c}_{1})=\ldots=\alpha(k)(\mathbb{P}^{c}_{k-1})\leq c$. Throughout the paper, we denote this value by $D(\alpha)$.
Elements of $\mathbb{P}^{c}$ will be used to record iterated degrees in bounded-degree akernels. For example, consider $\alpha\in\mathbb{P}^{0.8}$ such that
$\begin{split}\alpha(0)&=\star\;,\\ \alpha(1)&=0.8\cdot\mathsf{Dirac}(\star)\;,\\ \alpha(2)&=0.5\cdot\mathsf{Dirac}\left(0.1\cdot\mathsf{Dirac}(\star)\right)+0.3\cdot\mathsf{Dirac}\left(0.7\cdot\mathsf{Dirac}(\star)\right)\;,\\ \ldots\;.\end{split}$ (8)
If such an $\alpha$ encodes information about the iterated degrees in an akernel $W$ at a vertex $x$, then this means that
$\deg_{W}(x)=0.8,$ (9)
$\int_{y:\deg_{W}(y)=0.1}W(x,y)=0.5\;,\;\int_{y:\deg_{W}(y)=0.7}W(x,y)=0.3\;\mbox{,}\;\int_{y:\deg_{W}(y)\not\in\{0.1,0.7\}}W(x,y)=0.$ (10)
In particular, the condition $\alpha(k)=(p^{c}_{k-1,k})_{*}\alpha(k+1)$ in (6) expresses consistency of the iterated degrees. To give a specific example of violating this consistency condition, consider (10) but modify (9) to $\deg_{W}(x)=0.9$.
We now recall a certain operation of lifting elements of $\mathbb{P}^{c}$ to measures in $\mathcal{M}_{\leq c}(\mathbb{P}^{c})$ from [11]. Take the exemplar $\alpha\in\mathbb{P}^{1}$ from (8) to gain some intuition. This $\alpha$ is in a way expressing an object whose total mass is 0.8, and this total mass is split in the ratio $0.5:0.3$ between objects whose total mass is $0.1$ and $0.7$, respectively. We could equivalently encode such information into an object $\mu_{\alpha}\in\mathcal{M}_{\leq 1}(\mathbb{P}^{1})$ such that $\mu_{\alpha}(\mathbb{P}^{1})=0.8$, and $\mu_{\alpha}(D_{0.1})=0.5$ and $\mu_{\alpha}(D_{0.7})=0.3$, where for $r\geq 0$, $D_{r}:=\{\beta\in\mathbb{P}^{1}:D(\beta)=r\}$. The general definition is as follows. For $\alpha\in\mathbb{P}^{c}$ let $\mu_{\alpha}\in\mathcal{M}_{\leq c}(\mathbb{P}^{c})$ be the unique measure that satisfies
$(p_{k,\infty})_{*}\mu_{\alpha}=\alpha(k+1)$ (11)
for every $k\in\mathbb{N}$. The soundness of this definition follows from Kolmogorov’s extension theorem together with the consistency condition $\alpha(k)=(p^{c}_{k-1,k})_{*}\alpha(k+1)$ in (6).
Let $W$ be a bounded-degree akernel with $\operatorname{deg^{max}}(W)\leq c$. For $k=0,1,2\ldots$ we recursively define $i_{W,k}:X\to\mathbb{P}^{c}_{k}$ in the following way. Let $i_{W,0}\colon X\to\mathbb{P}^{c}_{0}=\{\star\}$ be the constant map. For $k\in\mathbb{N}_{0}$, let $i_{W,k+1}$ be constructed from $i_{W,k}$ so that
[…]
That is, the function $i_{W,1}$ carries information about degrees of vertices in the akernel $W$, the function $i_{W,2}$ carries information about a refined version of degrees such as ‘the degree of a vertex into the set of vertices whose degree lies in the interval $[0.4,0.5]$’, and so on. Let $i_{W}\colon X\to\mathbb{P}^{c}$ be defined by $i_{W}(x)(k)=i_{W,n}(x)(k)$. Analogously to the proof of Proposition 6.8 of [11], the map $i_{W}$ is well defined, i. e., we have $i_{W}(x)\in\mathbb{P}^{c}$ for every $x\in X$. Finally, let $\nu_{W}=(i_{W})_{*}\mu$ be the pushforward of $\mu$ via $i_{W}$. Obviously, we have $\nu_{W}\in\mathcal{M}_{=1}(\mathbb{P}^{c})$. We call $\nu_{W}$ the iterated degree measure of the bounded-degree akernel $W$.
Remark 2.2 (The bound on $c$ in the definition of $\mathbb{P}^{c}$).
Our main results assume some type of boundedness of the kernel or akernel in question. The mildest such assumption is $\operatorname{deg^{max}}(W)\leq c$. This assumption guarantees that the support of $\nu_{W}$ is contained in $\mathbb{P}^{c}$. To work with kernels without bounded degrees, we could easily repeat the construction and build a space $\mathbb{P}^{<\infty}$ using sets $L_{k+1}^{<\infty}:=\mathcal{M}_{<\infty}(\mathbb{P}_{k}^{<\infty})$. However, by extending from $\mathbb{P}^{c}$ to $\mathbb{P}^{<\infty}$ we lose compactness, a prominent feature needed for an application of the Stone–Weierstraß Theorem which in turn plays an important role in our proof of Theorem 1.3. A discussion of possible generalizations can be found in Section 7.
This is the only role the bound $c$ plays, and its value is fairly immaterial. This is reflected by the fact that we usually assume the value of $c$ implicitly fixed and remove it from main quantification of our lemmas.
Out of several equivalent definitions of fractional isomorphism from [11], we recall the one based on iterated degree measures.
Definition 2.3.
We say that two bounded-degree akernels $U$ and $W$ are fractionally isomorphic if $\nu_{U}=\nu_{W}$.
The last result in this section rectifies an omission from Section 1.2.
Lemma 2.4.
Suppose that $U$ and $W$ are two projectively fractionally isomorphic $L^{\infty}$-graphons. Then the constants $c_{U}$ and $c_{W}$ defined in (1) are equal.
Proof.
Suppose that $U$ is fractionally isomorphic to $tW$ for some $t>0$, i. e., $\nu_{U}=\nu_{tW}$. Looking at (1), we can replace $\deg_{W}(x)$ by $D(i_{W}(x))$. Also, $\|W\|_{1}=\int_{x}\deg_{W}(x)\mathsf{d}\mu(x)=\int D(\alpha)\mathsf{d}\nu_{W}(\alpha)$. Hence,
$c_{U}^{2}$ $=\frac{\int_{x}\deg_{U}(x)^{2}\mathsf{d}\mu(x)}{\|U\|_{1}^{2}}=\frac{\int D(\alpha)^{2}\mathsf{d}\nu_{U}(\alpha)}{\left(\int D(\alpha)\mathsf{d}\nu_{U}(\alpha)\right)^{2}}=\frac{\int D(\alpha)^{2}\mathsf{d}\nu_{tW}(\alpha)}{\left(\int D(\alpha)\mathsf{d}\nu_{tW}(\alpha)\right)^{2}}=\frac{\int_{x}\deg_{tW}(x)^{2}\mathsf{d}\mu(x)}{\|tW\|_{1}^{2}}$
$=\frac{\int_{x}t^{2}\deg_{W}(x)^{2}\mathsf{d}\mu(x)}{t^{2}\|W\|_{1}^{2}}=c_{W}^{2}.$
2.3 The Stone–Weierstraß Theorem
We recall the setting of the Stone–Weierstraß Theorem. Given a topological space $Z$, we write $C(Z,\mathbb{R})$ for the set of all continuous functions from $Z$ to $\mathbb{R}$. Recall that a family $\mathcal{E}\subset C(Z,\mathbb{R})$ is multiplicative if for every $f,g\in\mathcal{E}$ the function $h$ defined as $h(z)=f(z)g(z)$ satisfies $h\in\mathcal{E}$. We call a family $\mathcal{E}\subset C(Z,\mathbb{R})$ is multiplicative up to constants, if for every $f,g\in\mathcal{E}$ there exists an $a\in\mathbb{R}\setminus\{0\}$ such that the function $h$ defined as $h(z)=af(z)g(z)$ lies in $\mathcal{E}$. Given distinct elements $x,y\in Z$, the family $\mathcal{E}$ separates $x$ and $y$ if there exists $f\in\mathcal{E}$ such that $f(x)\neq f(y)$; the family $\mathcal{E}$ separates points if it separates every pair of distinct elements $x,y\in Z$.
Like in [11] we use a corollary of the Real Stone–Weierstraß Theorem, which can be found in [24, Theorem 7.32].
Corollary 2.5.
Let $K$ be a compact metric space and $c>0$ be a real number. Suppose that $\mathcal{E}\subset C(K,\mathbb{R})$ is a family of functions that is multiplicative up to constants, contains the constant-1 function, and separates points. Then for every distinct $\alpha,\beta\in\mathcal{M}_{\leq c}(K)$ there is $f\in\mathcal{E}$ such that
$\int_{K}f\mathsf{d}\alpha\neq\int_{K}f\mathsf{d}\beta\;.$
3 Proof of Theorem 1.3
For a kernel $W$, the offspring in the branching process $\mathfrak{X}_{W}$ are generated via a Poisson point process on $W$. Unfolding the definition of fractional isomorphism, Theorem 1.3 says that it is not the entire complexity of $W$ that determines $\mathfrak{X}_{W}$ but rather less complex information encoded in the iterated degrees.
In Section 3.1 we show that the fractional isomorphism type of a bounded-degree akernel $W$ determines the distribution of $\mathfrak{X}_{W}$ by constructing a process $\mathfrak{B}_{W}$ that depends only on the iterated degrees of vertices and not on the vertices themselves. In particular, this establishes the ()($\Leftarrow$)( ⇐ ) direction of Theorem 1.3.
In Section 3.2 we express the probabilities that $\mathfrak{B}_{W}$ (or rather, a related process $\mathfrak{B}(\alpha)$) generates a given tree. This is in particular used to establish, in Lemma 3.3, that certain families $\overline{\mathcal{F}_{n}}$ of functions on $\mathbb{P}_{n}^{c}$ which are related to these tree-generation probabilities are multiplicative up to constants. We also show in Lemma 3.5 that the functions in $\overline{\mathcal{F}_{n}}$ are continuous. In Section 3.3 we state Proposition 3.6, which asserts that $\overline{\mathcal{F}_{n}}$ also separates points. This puts us in a position where we can use the Stone–Weierstraß Theorem to show that for each pair of non-fractionally-isomorphic akernels $U$ and $W$ there is a rooted tree $T$ and $k\in\mathbb{N}_{0}$ such that $\mathbf{P}\left[(\mathfrak{X}_{U})_{\restriction k}\cong T\right]\neq\mathbf{P}\left[(\mathfrak{X}_{W})_{\restriction k}\cong T\right]$, hence establishing the ()($\Rightarrow$)( ⇒ ) direction of Theorem 1.3. We include the deferred proof of Proposition 3.6 in Section 3.4.
3.1 The ()($\Leftarrow$)( ⇐ ) direction of Theorem 1.3
For a bounded-degree akernel $W$, we introduce a branching process $\mathfrak{B}_{W}$ with type space $\mathbb{P}^{c}$. We show in Proposition 3.1 that this branching process is closely related to the branching process $\mathfrak{X}_{W}$.
Recall that for each $\beta\in\mathbb{P}^{c}$ we have a unique measure $\mu_{\beta}\in\mathcal{M}_{\leq c}(\mathbb{P}^{c})$ satisfying (11). Given any $\alpha\in\mathbb{P}^{c}$, the branching process $\mathfrak{B}(\alpha)$ is given as follows. The type space of $\mathfrak{B}(\alpha)$ is the set $\mathbb{P}^{c}$. The root has type $\alpha$. A particle of type $\beta\in\mathbb{P}^{c}$ has children distributed as a Poisson point process on $\mathbb{P}^{c}$ with intensity $\mu_{\beta}$. Given a bounded-degree akernel $W$, the branching process $\mathfrak{B}_{W}$ with type space $\mathbb{P}^{c}$ is given as follows. The root has type according to the distribution $\nu_{W}$ and a particle of type $\beta\in\mathbb{P}^{c}$ has children distributed as a Poisson point process on $\mathbb{P}^{c}$ with intensity $\mu_{\beta}$.
The following fact encapsulates the close connection between the branching processes $\mathfrak{B}_{W}$ and $\mathfrak{X}_{W}$ via the function $i_{W}$ defined in Section 2.2. In particular, Proposition 3.1(iii) establishes the ()($\Leftarrow$)( ⇐ ) direction of Theorem 1.3. For $x\in X$, we write $\mathfrak{X}_{W}(x)$ for the branching process $\mathfrak{X}_{W}$ conditioned to start at $x$.
Proposition 3.1.
  1. (i)
    Suppose that we are given a bounded-degree akernel $W$. For each $x\in X$, the branching processes $\mathfrak{X}_{W}(x)$ and $\mathfrak{B}(i_{W}(x))$ have the same distribution.
  2. (ii)
    Suppose that we are given a bounded-degree akernel $W$. The branching processes $\mathfrak{X}_{W}$ and $\mathfrak{B}_{W}$ have the same distribution.
  3. (iii)
    Suppose that we are given two fractionally isomorphic bounded-degree akernels $U$ and $W$. Then $\mathfrak{X}_{U}$ and $\mathfrak{X}_{W}$ have the same distribution.
Proof.
The implication (i)$\Rightarrow$(ii) is clear because $\nu_{W}$ is the pushforward of $\mu$ via $i_{W}$. The implication (ii)$\Rightarrow$(iii) is also clear because two fractionally isomorphic bounded-degree akernels $U$ and $W$ have the same iterated degree measure $\nu_{U}=\nu_{W}$ by definition. It remains to prove (i).
While we generally ignore type labels and compare branching processes via their distributions on type-unlabeled rooted trees, here we will retain the type labels. More specifically, we prove the desired statement by showing that we may couple the distributions of $\mathfrak{X}_{W}(x)$ and $\mathfrak{B}(i_{W}(x))$ on type-labeled rooted trees by using the natural transformation $x\mapsto i_{W}(x)$ on particle types.
We show by induction on $k$ that for every $k\in\mathbb{N}_{0}$ and $x\in X$, the $k$-balls $(\mathfrak{X}_{W}(x))_{\restriction k}$ and $(\mathfrak{B}(i_{W}(x)))_{\restriction k}$ have the same distribution on type-labeled rooted trees. This is trivial for $k=0$. For step $k\in\mathbb{N}$ observe that $(\mathfrak{X}_{W}(x))_{\restriction k}$ consists of the root of type $x$, children of types $x_{1},\ldots,x_{h}\in X$ following a Poisson point process on $X$ with intensity $\pi$ given by $\mathsf{d}\pi=W(x,\cdot)\mathsf{d}\mu$, and trees $(\mathfrak{X}_{W}(x_{1}))_{\restriction k-1}$, …, $(\mathfrak{X}_{W}(x_{h}))_{\restriction k-1}$ attached to them. On the other hand, $(\mathfrak{B}(i_{W}(x)))_{\restriction k}$ consists of the root of type $i_{W}(x)$, children of types $\alpha_{1},\ldots,\alpha_{\ell}\in\mathbb{P}^{c}$ following a Poisson point process on $\mathbb{P}^{c}$ with intensity $\mu_{\alpha}$, and trees $(\mathfrak{B}(\alpha_{1}))_{\restriction k-1}$, …, $(\mathfrak{B}(\alpha_{\ell}))_{\restriction k-1}$ attached to them. Now by the Mapping Theorem for point processes, $i_{W}(x_{1}),\ldots,i_{W}(x_{h})$ follows a Poisson point process on $\mathbb{P}^{c}$ with intensity $(i_{W})_{*}\pi$. By the definitions of pushforward measure, $i_{W}$ and $\mu_{i_{W}(x)}$ (see (13) and (11)), we have $(i_{W})_{*}\pi=\mu_{i_{W}(x)}$. Hence, the inductive step is completed by the inductive hypothesis. ∎
3.2 The probability of generating a tree
In this section we introduce notation that will allow us to express the probability that the first $k$ levels of $\mathfrak{B}(\alpha)$ yield a given tree. Constructions to this end are very similar to [11, Section 7.2] in which measures $\alpha\in\mathbb{P}^{1}$ were used to express the rooted homomorphism density of a rooted tree. While there is some resemblance between our branching processes and homomorphism densities of trees, the two concepts are not directly related.
For a finite rooted tree $T$ with root $r$, its height is given by $h(T):=\max(dist(r,v)\mid v\in V(T))$. For $n\in\mathbb{N}$ let $\mathcal{T}_{n}$ be the set of isomorphism classes of finite rooted trees of height at most $n$. We denote the $1$-vertex rooted tree by $\square$, i. e., $\mathcal{T}_{0}=\{\square\}$.
We introduce two basic operations for building a rooted tree from smaller ones. The first one plants a new root, while the second one merges several trees into one by identifying their roots.
  • Suppose that $T$ is a rooted tree with root $r$. Then let $T^{\uparrow}$ denote the tree consisting of the unrooted version of $T$ with a new root attached to the previous root. In particular, we have $v(T^{\uparrow})=v(T)+1$ and $h(T^{\uparrow})=h(T)+1$.
  • Suppose that $T_{1},\ldots,T_{\ell}$ are finite rooted trees. The rooted tree $T_{1}\oplus\ldots\oplus T_{\ell}$ is constructed by taking their disjoint union while identifying all the roots. In particular, we have $v(T_{1}\oplus\ldots\oplus T_{\ell})=v(T_{1})+\cdots+v(T_{\ell})-\ell+1$ and $h(T_{1}\oplus\ldots\oplus T_{\ell})=\max\{h(T_{1}),\ldots,h(T_{\ell})\}$.
For each finite rooted tree $T$ and each $k\in\mathbb{N}_{0}$, we define two functions $f_{T,k},\overline{f_{T,k}}\colon\mathbb{P}_{k}^{c}\to\mathbb{R}$. Lemma 3.4 connects our algebraic Definition 3.2 of $f_{T,n}(p_{n,\infty}(\alpha))$ to the branching process $\mathfrak{B}(\alpha)$. The way we define $\overline{f_{T,n}}(\beta)$ (in Definition 3.2(iv)) based on $f_{T,n}(\beta)$ is quite bland, yet convenient for the application of the Stone–Weierstraß Theorem.
Suppose that $F$ is a finite rooted tree with root $r$. Let $\mathcal{L}_{F}=\{\ell_{1},\ldots\ell_{t}\}$ be the multiset of multiplicities of isomorphism types of connected components of $F-r$. For example, $\mathcal{L}_{F}=\{1,3,3\}$ means that $F-r$ has 7 connected components in total, comprising one triple of mutually isomorphic connected components, another triple of mutually isomorphic connected components, and a further unique connected component. Note that $\sum\mathcal{L}_{F}=\deg_{F}(r)$. Define a constant $e_{F}$ using elements in the multiset $\mathcal{L}_{F}$,
$e_{F}:=\prod_{\ell\in\mathcal{L}_{F}}\frac{1}{\ell!}\;.$ (14)
Definition 3.2.
  1. (i)
    For each $k\in\mathbb{N}_{0}$ and each $\beta\in\mathbb{P}_{k}^{c}$, we have
    $f_{\square,k}(\beta)=\begin{cases}1&\text{ if $k=0$,}\\ \exp(-D(\beta))&\text{ if $k>0$.}\end{cases}$
  2. (ii)
    For a finite rooted tree $T$ of height at most $k\in\mathbb{N}_{0}$ and $\beta\in\mathbb{P}_{k+1}^{c}$ define
    $f_{T^{\uparrow},k+1}(\beta)=\exp(-D(\beta))\cdot\int_{\sigma\in\mathbb{P}_{k}^{c}}f_{T,k}(\sigma)\mathsf{d}\big{(}\beta(k+1)\big{)}(\sigma)\;.$
  3. (iii)
    For finite rooted trees $T_{1},\ldots,T_{\ell}$ of positive heights not exceeding $k\in\mathbb{N}$ and whose roots have degree exactly 1, and for $\beta\in\mathbb{P}_{k}^{c}$,
    $f_{T_{1}\oplus\ldots\oplus T_{\ell},k}(\beta)=e_{T_{1}\oplus\ldots\oplus T_{\ell}}\cdot\exp(D(\beta))^{\ell-1}\prod_{i=1}^{\ell}f_{T_{i},k}(\beta)\;.$
    Note that this definition is consistent for $\ell=1$, that is, it says $f_{T_{1},k}(\beta)=1\cdot\exp(D(\beta))^{0}f_{T_{1},k}(\beta)$.
  4. (iv)
    For each finite rooted tree $T$ and each $k\in\mathbb{N}_{0}$, we define $\overline{f_{T,k}}(\beta)=\exp(D(\beta))f_{T,k}(\beta)$. In particular, we have $\overline{f_{\square,1}}(\beta)=1$.
Every rooted tree can be decomposed into a collection of subtrees rooted at the children of the original root, and we can perform this decomposition recursively. Hence, the function $f_{T,k}$ is given for every $k\in\mathbb{N}_{0}$ and for every finite tree $T$ with $h(T)\leq k$. Let $\mathcal{F}_{n}:=\{f_{T,n}\mid T\in\mathcal{T}_{n}\}$ and $\overline{\mathcal{F}_{n}}:=\{\overline{f_{T,n}}\mid T\in\mathcal{T}_{n}\}$.
The next lemma is needed to apply Corollary 2.5.
Lemma 3.3.
For each $n\in\mathbb{N}$, the set $\overline{\mathcal{F}_{n}}$ is multiplicative up to constants.
Proof.
Let $T$ and $F$ be finite rooted trees of height at most $n$. We shall show that $\overline{f_{T,n}}\cdot\overline{f_{F,n}}=\frac{e_{T}e_{F}}{e_{T\oplus F}}\cdot\overline{f_{T\oplus F,n}}$. As $h(T\oplus F)=\max(h(T),h(F))$, the function $\overline{f_{T\oplus F,n}}$ is contained in $\overline{\mathcal{F}_{n}}$. This will show multiplicativity up to constants.
Let $T_{1},\ldots,T_{\ell}$ be the connected components of $T$ after the removal of the root. Let $F_{1},\ldots,F_{t}$ be the connected components of $F$ after the removal of the root. Note that the connected components of $T\oplus F$ after the removal of the root are $T_{1},\ldots,T_{\ell},F_{1},\ldots,F_{t}$. For each $\alpha\in\mathbb{P}_{n}^{c},$ we have
$\overline{f_{T,n}}(\alpha)\cdot\overline{f_{F,n}}(\alpha)$ $=\exp(D(\alpha))^{2}f_{T,n}(\alpha)\cdot f_{F,n}(\alpha)$
$=\exp(D(\alpha))^{2}\cdot e_{T}\cdot\exp(D(\alpha))^{\ell-1}\prod_{i=1}^{\ell}f_{T_{i}^{\uparrow},n}(\alpha)\cdot e_{F}\cdot\exp(D(\alpha))^{t-1}\prod_{i=1}^{t}f_{F_{i}^{\uparrow},n}(\alpha)$
$=\frac{e_{T}e_{F}}{e_{T\oplus F}}\cdot\exp(D(\alpha))\cdot f_{T\oplus F,n}(\alpha)=\frac{e_{T}e_{F}}{e_{T\oplus F}}\cdot\overline{f_{T\oplus F,n}}(\alpha)\;.\qed$
The next lemma connects the above analytic definition of functions $f_{T,n}(\cdot)$ with branching processes $\mathfrak{B}(\cdot)$.
Lemma 3.4.
Given $k\in\mathbb{N}_{0}$, a rooted tree $T$ of height at most $k$, and $\alpha\in\mathbb{P}^{c}$, the quantity $f_{T,k}(p^{c}_{k,\infty}(\alpha))$ is the probability that the first $k$ levels of $\mathfrak{B}(\alpha)$ are isomorphic to $T$.
Proof.
The claim is trivial for $k=0$. Indeed, the zeroth level of $\mathfrak{B}(\alpha)$ is always isomorphic to $T=\square$, the only tree of height at most 00, and we have $f_{\square,0}(p^{c}_{0,\infty}(\alpha))=1$.
For $k>0$ we proceed by induction on $k$. Let $r$ be the root of $T$. Write $s=\deg_{T}(r)$. Let $\mathcal{L}_{T}=\{\ell_{1},\ldots\ell_{t}\}$ be the multiset of multiplicities of isomorphism types of connected components of $T-r$. Let $T_{1},\ldots,T_{t}$ be the isomorphism types of these connected components of $T-r$ corresponding to these multiplicities.
First, let us argue when $D(\alpha)=0$. Indeed then on the one hand, $\mu_{\alpha}$ is the trivial measure and $\mathfrak{B}(\alpha)$ is almost surely just the root. Let us now turn to evaluating $f_{T,k}(p^{c}_{k,\infty}(\alpha))$. If $T=\square$ then Definition 3.2(i) gives that $f_{T,k}(p^{c}_{k,\infty}(\alpha))=\exp(-D(\alpha))=1$. If $T\neq\square$, then $r$ has at least one nontrivial subtree attached to it. Using the product formula of Definition 3.2(iii), in order to show that $f_{T,k}(p^{c}_{k,\infty}(\alpha))=0$, it suffices to see nullification on this one subtree. This can be seen from the integral formula in Definition 3.2(ii) as the measure $\alpha(k)$ is trivial.
Thus, in the following we can assume that $D(\alpha)>0$. Generate the first generation of $\mathfrak{B}(\alpha)$. If the first $k$ levels of $\mathfrak{B}(\alpha)$ are isomorphic to $T$ then there are $s$ many offspring generated in the first generation. We shall assume that this occurred. Let $[s]$ be an enumeration of the offspring of the root considered in a uniformly random order. For a given partition $\mathcal{U}\in\binom{[s]}{\mathcal{L}_{F}}$, let $E_{\mathcal{U}}$ be the event for each $i\in[t]$ the $\ell_{i}$ many elements of $[s]$ corresponding to the $i$th type of $\mathcal{U}$ generate a tree isomorphic to $T_{i}$. The events $E_{\mathcal{U}}$ are disjoint, their union is the event that the first $k$ levels of $\mathfrak{B}(\alpha)$ are isomorphic to $T$, and (because the ordering is random) they all occur with the same probability. In particular, we can focus on the partition $\mathcal{U}^{*}=(\{1,\ldots,\ell_{1}\},\{\ell_{1}+1,\ldots,\ell_{1}+\ell_{2}\},\ldots,\{\ell_{1}+\ell_{2}+\ldots+\ell_{t-1}+1,\ldots,s\})$. This gives that
$\mathbf{P}\left[\mathfrak{B}(\alpha)_{\restriction k}\cong T\right]=\binom{s}{\mathcal{L}_{F}}\cdot\frac{D(\alpha)^{s}\exp(-D(\alpha))}{s!}\cdot\prod_{i=1}^{t}p_{i}^{\ell_{i}}\;,$ (15)
where $p_{i}$ is the probability that $T_{i}$ is isomorphic to the first $k-1$ level of the branching process $\mathfrak{B}(\beta)$, where $\beta\in\mathbb{P}_{k-1}^{c}$ is chosen at random according to the probability distribution $\nicefrac{{\alpha(k)}}{{D(\alpha)}}$. That is, using the induction hypothesis, we have
$p_{i}=\int_{\beta\in\mathbb{P}_{k-1}^{c}}f_{T_{i},k-1}(\beta)\;\mathsf{d}\left(\nicefrac{{\alpha(k)}}{{D(\alpha)}}\right)(\beta)\;.$ (16)
Putting (15) and (16) together we have
$\mathbf{P}\left[\mathfrak{B}(\alpha)_{\restriction k}\cong T\right]=\frac{D(\alpha)^{s}\exp(-D(\alpha))}{\prod_{i=1}^{t}\ell_{i}!}\cdot\prod_{i=1}^{t}p_{i}^{\ell_{i}}=\frac{\exp(-D(\alpha))}{\prod_{i=1}^{t}\ell_{i}!}\cdot\prod_{i=1}^{t}\left(\int_{\sigma\in\mathbb{P}_{k-1}^{c}}f_{T_{i},k-1}(\sigma)\;\mathsf{d}(\alpha(k))(\sigma)\right)^{\ell_{i}}\;.$ (17)
We now shift to expanding $f_{T,k}(\alpha)$. We first use Definition 3.2(iii) and then Definition 3.2(ii).
$f_{T,k}(p^{c}_{k,\infty}(\alpha))$ $=\frac{1}{\prod_{i=1}^{t}\ell_{i}!}\cdot\exp(D(\alpha))^{s-1}\prod_{i=1}^{t}\left(f_{T_{i}^{\uparrow},k}(p^{c}_{k,\infty}(\alpha))\right)^{\ell_{i}}$ (18)
$=\frac{1}{\prod_{i=1}^{t}\ell_{i}!}\cdot\exp(D(\alpha))^{s-1}\cdot\exp(-D(\alpha))^{s}\cdot\prod_{i=1}^{t}\left(\int_{\sigma\in\mathbb{P}_{k}^{c}}f_{T_{i},k-1}(\sigma)\mathsf{d}\big{(}\alpha(k)\big{)}(\sigma)\right)^{\ell_{i}}$
$=\frac{\exp(-D(\alpha))}{\prod_{i=1}^{t}\ell_{i}!}\cdot\prod_{i=1}^{t}\left(\int_{\sigma\in\mathbb{P}_{k}^{c}}f_{T_{i},k-1}(\sigma)\mathsf{d}\big{(}\alpha(k)\big{)}(\sigma)\right)^{\ell_{i}}\;.$ (19)
This finishes the proof. ∎
The Stone–Weierstraß Theorem works in the setting of continuous functions. The next lemma shows that this is indeed the case for the functions $f_{T,n}$.
Lemma 3.5.
Given $k\in\mathbb{N}_{0}$ and a rooted tree $T$ of height at most $k$, the functions $f_{T,k}:\mathbb{P}_{k}^{c}\to\mathbb{R}$ and $\overline{f_{T,k}}:\mathbb{P}_{k}^{c}\to\mathbb{R}$ are continuous.
Proof.
Let us first think about the topology on $\mathbb{P}^{c}_{k}$. As a matter of fact, this is the only place in the paper where we need to study the topology on $\mathbb{P}^{c}_{k}$. The only topological fact about $\mathbb{P}^{c}_{k}$ used at other places is that of compactness. Recall that the topologies on $L_{i}^{c}$ are metrizable, and hence so is the product topology on $\mathbb{P}^{c}_{k}$. Thus the topology is characterized by convergent sequences.
Claim.
Let $k\in\mathbb{N}$ be arbitrary. Suppose that $\alpha_{1},\alpha_{2},\ldots\in\mathbb{P}^{c}_{k}$ converge in $\mathbb{P}^{c}_{k}$ to $\alpha\in\mathbb{P}^{c}_{k}$. Then for every bounded continuous function $f:\mathbb{P}^{c}_{k-1}\rightarrow\mathbb{R}$ we have
$\lim_{n\to\infty}\int_{\sigma}f(\sigma)\mathsf{d}\big{(}\alpha_{n}(k)\big{)}(\sigma)=\int_{\sigma}f(\sigma)\mathsf{d}\big{(}\alpha(k)\big{)}(\sigma)\;.$ (20)
Proof.
Indeed, since $\alpha_{1},\alpha_{2},\ldots$ converge in $\mathbb{P}^{c}_{k}$ to $\alpha$, then in particular $\alpha_{1}(k),\alpha_{2}(k),\ldots$ converge in $L^{c}_{k}$ to $\alpha(k)$. By (5), we have that $L^{c}_{k}=\mathcal{M}_{\leq c}(\mathbb{P}_{k-1}^{c})$ is equipped by the weak topology, of which (20) is the defining property (c.f. (2)). ∎
Observe that the total mass function $D:\mathbb{P}_{k}^{c}\to[0,c]$ is continuous. Hence, it is enough to focus on functions $f_{T,k}:\mathbb{P}_{k}^{c}\to\mathbb{R}$. For each $k$, the function $f_{\square,k}$ is continuous by its definition given in Definition 3.2(i). For more complicated trees, we proceed by induction on $k$. Reflecting Definition 3.2, we distinguish whether the root of $T$ has degree 1 or more. In the former case we can write $T=F^{\uparrow}$. We apply Definition 3.2(ii),
$f_{T,k}(\beta)=\exp(-D(\beta))\cdot\int_{\sigma\in\mathbb{P}_{k-1}^{c}}f_{F,k-1}(\sigma)\mathsf{d}\big{(}\beta(k)\big{)}(\sigma)\;.$
The term $\exp(-D(\beta))$ is continuous in $\beta$ as we noted above. The term $\int_{\sigma\in\mathbb{P}_{k-1}^{c}}f_{F,k-1}(\sigma)\mathsf{d}\big{(}\beta(k)\big{)}(\sigma)$ is continuous in $\beta$ by Claim and the inductive assumption about the continuity of the function $f_{F,k-1}$.
If the root of $T$ has degree more than 1, then we use Definition 3.2(iii). It tells us that we can write $f_{T,k}$ as a product of a constant, a power of $\exp(-D(\beta))$, and a product of functions $f_{T_{i},k}$ where the degree of the root in each $T_{i}$ is 1. In particular, the functions $f_{T_{i},k}$ are continuous as per the previous case. Hence, $f_{T,k}$ is continuous. ∎
3.3 Finding a tree separating iterated degree sequences
The following proposition is the key step to proving the ()($\Rightarrow$)( ⇒ ) direction of Theorem 1.3. Its proof is given in Section 3.4.
Proposition 3.6.
For each $n\in\mathbb{N}_{0}$ the sets $\mathcal{F}_{n}$ and $\overline{\mathcal{F}_{n}}$ both separate points of $\mathbb{P}_{n}^{c}$.
To finish the proof of Theorem 1.3, the following concept will be useful. Given $c,d>0$, $n\in\mathbb{N}_{\infty}$ and a measure $\nu\in\mathcal{M}_{\leq d}(\mathbb{P}_{n}^{c})$, its exponential tilting is a measure $\widehat{\nu}$ on $\mathbb{P}_{n}^{c}$ defined for each Borel $S\subset\mathbb{P}_{n}^{c}$ by
$\widehat{\nu}(S):=\int_{\alpha\in S}\exp(-D(\alpha))\mathsf{d}\nu(\alpha)\;.$
The next lemma asserts that exponential tilting is injective on $\mathcal{M}_{\leq d}(\mathbb{P}_{n}^{c})$.
Lemma 3.7.
Suppose that $c,d>0$, $n\in\mathbb{N}_{\infty}$ and that $\nu_{1},\nu_{2}\in\mathcal{M}_{\leq d}(\mathbb{P}_{n}^{c})$ are two different measures. Then their exponential tiltings $\widehat{\nu_{1}}$ and $\widehat{\nu_{2}}$ are different measures in $\mathcal{M}_{\leq d}(\mathbb{P}_{n}^{c})$.
Proof.
The standard rules for Radon–Nikodym derivatives imply that exponential tilting is invertible, that is, for $i=1,2$ we have $\nu_{i}(S):=\int_{\alpha\in S}\exp(D(\alpha))\mathsf{d}\widehat{\nu_{i}}(\alpha)$ for each measurable subset $S\subset\mathbb{P}_{n}^{c}$. Since $\nu_{1}$ and $\nu_{2}$ are different measures, this implies that $\widehat{\nu_{1}}$ and $\widehat{\nu_{2}}$ are different. Furthermore, for $i=1,2$ the fact that $\exp(-D(\alpha))\leq 1$ for each $\alpha\in\mathbb{P}_{n}^{c}$ yields
$\widehat{\nu_{i}}(\mathbb{P}_{n}^{c})=\int_{\alpha\in\mathbb{P}_{n}^{c}}\exp(-D(\alpha))\mathsf{d}\nu_{i}(\alpha)\leq\int_{\alpha\in\mathbb{P}_{n}^{c}}1\mathsf{d}\nu_{i}(\alpha)=\nu_{i}(\mathbb{P}_{n}^{c})\leq d\;,$
so we have $\widehat{\nu_{i}}\in\mathcal{M}_{\leq d}(\mathbb{P}_{n}^{c})$ as required. ∎
Proof of the ()($\Rightarrow$)( ⇒ ) direction of Theorem 1.3.
Suppose that $U$ and $W$ are not fractionally isomorphic, that is, we have $\nu_{U}\neq\nu_{W}$. Thus, there is an $n\in\mathbb{N}$ such that $\nu_{U,n}\neq\nu_{W,n}$, where $\nu_{W,n}:=(p_{n,\infty}^{c})_{*}\nu_{W}$. Fix that $n$ and let $\widehat{\nu_{U,n}}$ and $\widehat{\nu_{W,n}}$ be exponential tiltings of $\nu_{U,n}$ and $\nu_{W,n}$ respectively. By Lemma 3.7 we have $\widehat{\nu_{U,n}}\neq\widehat{\nu_{W,n}}$.
The family $\overline{\mathcal{F}_{n}}$ consists of continuous functions (Lemma 3.5), is multiplicative up to constants (by Lemma 3.3), contains the constant-1 function (by Definition 3.2(iv)), and separates points of $\mathbb{P}_{n}^{c}$ (by Proposition 3.6). Therefore, the Stone–Weierstraß Theorem (Corollary 2.5) tells us that there exists a finite rooted tree $T$ of height at most $n$ such that
$\int_{\alpha\in\mathbb{P}_{n}^{c}}\overline{f_{T,n}}(\alpha)\mathsf{d}\widehat{\nu_{U,n}}(\alpha)\neq\int_{\alpha\in\mathbb{P}_{n}^{c}}\overline{f_{T,n}}(\alpha)\mathsf{d}\widehat{\nu_{W,n}}(\alpha)\;.$ (21)
By unpacking the definitions of $\overline{f_{T,n}}$ and $\widehat{\nu_{U,n}}$, and then applying Lemma 3.4 with $f_{T,n}$, we obtain
$\mathbf{P}[(\mathfrak{X}_{U})_{\restriction n}\cong T]=\int_{\alpha\in\mathbb{P}^{c}}f_{T,n}(p^{c}_{n,\infty}(\alpha))\mathsf{d}\nu_{U}(\alpha)=\int_{\sigma\in\mathbb{P}^{c}_{n}}f_{T,n}(\sigma)\mathsf{d}\nu_{U,n}(\sigma)=\int_{\sigma\in\mathbb{P}_{n}^{c}}\overline{f_{T,n}}(\sigma)\mathsf{d}\widehat{\nu_{U,n}}(\sigma)\;.$
Analogously, we have $\mathbf{P}[(\mathfrak{X}_{W})_{\restriction k}\cong T]=\int_{\alpha\in\mathbb{P}_{n}^{c}}\overline{f_{T,n}}(\sigma)\mathsf{d}\widehat{\nu_{W,n}}(\sigma)$. Hence, by (21) we have $\mathbf{P}[(\mathfrak{X}_{U})_{\restriction k}\cong T]\neq\mathbf{P}[(\mathfrak{X}_{W})_{\restriction k}\cong T]$, completing the proof. ∎
3.4 Proof of Proposition 3.6
We shall prove by induction on $n\in\mathbb{N}_{0}$ that $\mathcal{F}_{n}$ and $\overline{\mathcal{F}_{n}}$ both separate any pair $\alpha\neq\beta\in\mathbb{P}_{n}^{c}$. By consistency, $\alpha\neq\beta$ implies $\alpha(n)\neq\beta(n)$.
The statement is trivial for $n=0$ because $\alpha(0)=\beta(0)=\mathbb{P}_{0}=\{\star\}$ has only one element. Now suppose that the claim has been shown for some $n\in\mathbb{N}_{0}$. Let $\alpha\neq\beta\in\mathbb{P}_{n+1}^{c}$. Let $\alpha^{\prime}$ and $\beta^{\prime}$ be the exponential tiltings of $\alpha(n+1)$ and $\beta(n+1)$ respectively. By Lemma 3.7, $\alpha^{\prime}$ and $\beta^{\prime}$ are different measures in $\mathcal{M}_{\leq c}(\mathbb{P}_{n}^{c})$.
By the inductive hypothesis, $\overline{\mathcal{F}_{n}}$ separates elements of $\mathbb{P}_{n}^{c}$. By Lemma 3.3, $\overline{\mathcal{F}_{n}}$ is multiplicative up to constants. Clearly $\overline{\mathcal{F}_{n}}$ contains the constant-1 function. Hence, the Stone–Weierstraß Theorem (Corollary 2.5) implies that there is a $T\in\mathcal{T}_{n}$ such that
$\int_{\mathbb{P}_{n}^{c}}\overline{f_{T,n}}(\gamma)\mathsf{d}\alpha^{\prime}(\gamma)\neq\int_{\mathbb{P}_{n}^{c}}\overline{f_{T,n}}(\gamma)\mathsf{d}\beta^{\prime}(\gamma)\;.$
We have $\overline{f_{T,n}}(\gamma)=\exp(D(\gamma))f_{T,n}(\gamma)$ by Definition 3.2(iv). Together with the definitions of $\alpha^{\prime}$ and $\beta^{\prime}$, this yields
$\int_{\mathbb{P}_{n}^{c}}f_{T,n}(\gamma)\exp(D(\gamma))\exp(-D(\gamma))\mathsf{d}\alpha(n+1)\neq\int_{\mathbb{P}_{n}^{c}}f_{T,n}(\gamma)\exp(D(\gamma))\exp(-D(\gamma))\mathsf{d}\beta(n+1),$
which cancels to
[…]
Thus, $\overline{\mathcal{F}_{n+1}}$ separates $\alpha$ and $\beta$.
To complete the inductive step, it remains to show that $\mathcal{F}_{n+1}$ also separates $\alpha$ and $\beta$. We have from before a rooted tree $T\in\mathcal{T}_{n}$ such that $\overline{f_{T^{\uparrow},n+1}}(\alpha)\neq\overline{f_{T^{\uparrow},n+1}}(\beta)$. We now distinguish two cases depending on whether $f_{T^{\uparrow},n+1}(\alpha)\neq f_{T^{\uparrow},n+1}(\beta)$ or $f_{T^{\uparrow},n+1}(\alpha)=f_{T^{\uparrow},n+1}(\beta)$. Obviously, $\mathcal{F}_{n+1}$ separates $\alpha$ and $\beta$ in the former case. So, let us assume the latter case. For all $\gamma\in\mathbb{P}^{c}$ we have
[…]
Using this for $\gamma=\alpha,\beta$, it follows that $f_{T^{\uparrow}\oplus T^{\uparrow},n+1}(\alpha)\neq f_{T^{\uparrow}\oplus T^{\uparrow},n+1}(\beta)$. This shows that $\mathcal{F}_{n+1}$ separates $\alpha$ and $\beta$, completing the proof.
4 More tools for proof of Theorem 1.7
4.1 Random walks on general state spaces
One of the key techniques in our proof is random walks on graphons. As far as we could find, the only work dealing with this topic is [22] which focuses on linking random walks on dense graphs to those on their limiting graphon. In this paper, we do not need this link and study just the graphon setting. To this end, standard theory of discrete-time Markov chains on general probability spaces suffices. We use [21] as our reference.
We begin with some definitions. Let $(X,\mathcal{B})$ be a standard Borel space. A function $P\colon X\times\mathcal{B}\to[0,1]$ is a transition probability kernel if for each $x\in X$ the measure $P_{x}$ on $\mathcal{B}$ given by $P_{x}(A)=P(x,A)$ is a probability measure and for each $A\in\mathcal{B}$ the function $f_{A}\colon X\to[0,1]$ given by $f_{A}(x)=P(x,A)$ is a measurable function on $X$. A time-homogeneous Markov chain with transition probability kernel $P$ and initial distribution $\chi$ is a stochastic process $\Phi=\{\Phi(n)\}_{n\in\mathbb{N}_{0}}$ defined on $\left(\prod_{i\in\mathbb{N}_{0}}X,\bigotimes_{i\in\mathbb{N}_{0}}\mathcal{B}\right)$ whose finite dimensional distributions are such that for all $n\in\mathbb{N}$ and measurable $A_{0},\dots,A_{n}\subset X$ we have
$\mathbf{P}\left[(\Phi(0)\in A_{0},\dots,\Phi(n)\in A_{n}\right]=\int_{y_{0}\in A_{0}}\cdots\int_{y_{n-1}\in A_{n-1}}P(y_{n-1},A_{n})\ \mathsf{d}P_{y_{n-2}}(y_{n-1})\dots\mathsf{d}P_{y_{0}}(y_{1})\ \mathsf{d}\chi(y_{0})\;.$
Given a transition probability kernel $P$, we recursively define the corresponding $n$-step transition probability kernel $P^{n}\colon X\times\mathcal{B}\to[0,\infty)$ as follows. Set $P^{0}(x,A)=\mathbbm{1}_{A}(x)$ for all $x\in X$ and $A\in\mathcal{B}$. For $n\in\mathbb{N}$ we recursively define
$P^{n}(x,A)=\int_{y\in X}P^{n-1}(y,A)\mathsf{d}P_{x}(y)$ (23)
for all $x\in X$ and $A\in\mathcal{B}$. We write $P^{n}_{x}$ for the probability measure on $\mathcal{B}$ given by $P^{n}_{x}(A):=P^{n}(x,A)$.
A key plank of Markov chain convergence is the property of irreducibility. Given a probability measure $\phi$ on $\mathcal{B}$, a Markov chain $\Phi$ is $\phi$-irreducible if for every $x\in X$ and every set $A\in\mathcal{B}$ with $\phi(A)>0$ there is a positive probability that $\Phi$ started at $x$ satisfies $\Phi(n)\in A$ for some $n\in\mathbb{N}$. We call $\phi$ an irreducibility measure for $\Phi$. The property of $\phi$-irreducibility is somewhat weak and fails to fully represent the usual notion of irreducibility for Markov chains on discrete state spaces, so we give a further definition to fully capture the range of behaviour of Markov chains. We say that a probability measure $\psi$ on $\mathcal{B}$ is a full irreducibility measure for $\Phi$ if there is an irreducibility probability measure $\phi$ for $\Phi$ such that, writing $P$ for the transition probability kernel of $\Phi$, for all $A\in\mathcal{B}$ we have
$\psi(A):=\int_{X}\sum_{j=1}^{\infty}2^{-j-1}P^{j}(y,A)\mathsf{d}\phi(y)\;.$
The concept of a full irreducibility probability measure is a cornerstone of the theory of Markov chains on general state spaces and provides the right setting for the study of Markov chain convergence. Indeed, it is shown in [21, Proposition 4.2.2] that full irreducibility probability measures have some additional properties that fully capture the range of behaviour of Markov chains. That said, we need full irreducibility probability measures solely to enable the application of Theorem 4.1 and will not directly utilize or mention any of the aforementioned properties. A more detailed discussion can be found in [21, Section 4.2].
A probability measure $\pi$ on $\mathcal{B}$ is invariant if for all $A\in\mathcal{B}$ we have $\pi(A)=\int P(x,A)\mathsf{d}\pi(x)$. A chain $\Phi$ is positive recurrent if it has a full irreducibility probability measure and an invariant probability measure.
Let $\Phi$ be a Markov chain with transition probability kernel $P$ and full irreducibility probability measure $\psi$, and let $p\in\mathbb{N}$. A $p$-cycle for $\Phi$ is a collection of disjoint sets $D_{p+1}=D_{1},\dots,D_{p}\in\mathcal{B}$ such that
  • for all $i\in[p]$ and all $x\in D_{i}$ we have $P(x,D_{i+1})=1$, and
  • the set $N=X\setminus(\bigcup_{i\in[p]}D_{i})$ satisfies $\psi(N)=0$.
The period of $\Phi$ is the largest $p\in\mathbb{N}$ for which $\Phi$ has a $p$-cycle. It follows from Theorems 5.2.2 and 5.4.4 in [21] that there exists $p\in\mathbb{N}$ such that $\Phi$ has a $p$-cycle and any $q\in\mathbb{N}$ for which $\Phi$ has a $q$-cycle must be a divisor of $p$; in particular, the period of $\Phi$ is well-defined.
A signed measure on $\mathcal{B}$ is a countably additive set function $\eta\colon\mathcal{B}\to(-\infty,\infty)$; in particular, this extends the usual notion of a measure by permitting negative values. The total variation norm of a signed measure $\eta$ on $\mathcal{B}$ is defined by
$\|\eta\|_{TV}=\sup_{f:X\rightarrow[-1,1]}\left|\int f(x)\mathsf{d}\eta(x)\right|=\sup_{A\in\mathcal{B}}\eta(A)-\inf_{A\in\mathcal{B}}\eta(A)\;.$
The main result from the theory of Markov chains we will need is the result about convergence of a Markov chain to the stationary distribution. Recall that in the more familiar setting of Markov chains on a finite or countable state space, in the aperiodic and irreducible case, the distribution of a random walk after $n$ steps converges to a uniquely defined stationary distribution in the total variation distance as $n\to\infty$, independently of the initial distribution. The following is a version of the convergence result for Markov chains on a general state space which handles periodicity by taking an average over each periodic interval.
Theorem 4.1 (Theorem 13.3.4(ii) in [21]).
Suppose that $\Phi$ is a positive recurrent Markov chain with period $p\in\mathbb{N}$ with an invariant probability measure $\pi$. Then for every initial probability distribution $\chi$ which is absolutely continuous with respect to $\pi$ we have
$\left\|\int\frac{1}{p}\sum_{r=0}^{p-1}P^{np+r}(x,\cdot)\ \mathsf{d}\chi(x)-\pi(\cdot)\right\|_{TV}\to 0\textrm{ as }n\to\infty.$
4.2 Markov renormalization $W^{\dagger}$
In this section, we introduce a way to transform a kernel $W$ into an akernel $W^{\dagger}$, which we call the Markov renormalization. The Markov renormalization will play a key role in the proof of Theorem 1.7. The defining formula comes from the density appearing in Definition 1.5(ii).
Definition 4.2.
Let $W$ be an $L^{\infty}$-kernel. The Markov renormalization of $W$, denoted $W^{\dagger}$, is an akernel $W^{\dagger}:X^{2}\rightarrow[0,\infty)$ defined at $(x,y)\in X^{2}$ by
$W^{\dagger}(x,y):=\frac{W(x,y)}{\deg_{W}(y)}\quad\text{with the convention $\frac{0}{0}=0$}\;.$
The reason why we call $W^{\dagger}$ the Markov renormalization is that for each $y\in X$ we have $\int_{x\in X}W^{\dagger}(x,y)=1$. Thus we can interpret $W^{\dagger}(\cdot,y)$ as transition densities from state $y$ of a Markov chain on $X$.
Lemma 4.3.
Suppose that $W$ is an $L^{\infty}$-kernel with minimum degree $\delta>0$. Then $W^{\dagger}$ is an $L^{\infty}$-akernel with $\|W^{\dagger}\|_{\infty}\leq\frac{\|W\|_{\infty}}{\delta}$ and minimum degree at least $\frac{\delta}{\|W\|_{\infty}}$.
Proof.
The quantity $\|W\|_{\infty}$ is an upper bound on the maximum degree of $W$, so for $\mu$-almost every $x\in X$ we have
$\deg_{W^{\dagger}}(x)=\int_{y}\frac{W(x,y)}{\deg_{W}(y)}\mathsf{d}\mu(y)\geq\int_{y}\frac{W(x,y)}{\|W\|_{\infty}}\mathsf{d}\mu(y)\geq\frac{\delta}{\|W\|_{\infty}}\;.$
This yields the desired lower bound on the minimum degree. The upper bound on $\|W^{\dagger}\|_{\infty}$ follows from the observation that we have $W^{\dagger}(x,y)=0$ or $W^{\dagger}(x,y)=\frac{W(x,y)}{\deg_{W}(y)}\leq\frac{\|W\|_{\infty}}{\delta}$ for $(\mu\times\mu)$-almost every $(x,y)\in X\times X$. ∎
The following fact follows directly by comparing the definitions of $\mathfrak{U}^{-}_{W}$ and $\mathfrak{X}_{W^{\dagger}}$.
Fact 4.4.
Suppose that $W$ is an $L^{\infty}$-kernel with positive minimum degree. Then the branching process $\mathfrak{U}^{-}_{W}$ defined in Figure 1 has the same distribution as $\mathfrak{X}_{W^{\dagger}}$.
4.2.1 Extinction
The main result of this section is Lemma 4.7, which asserts that for each kernel $W$ the branching process $\mathfrak{X}_{W^{\dagger}}$ goes extinct almost surely.
Lemma 4.5.
Given $\rho>0$ there exists a number $g_{\rho}>0$ such that the following holds. Suppose that $W$ is a kernel satisfying $\frac{\operatorname{deg^{min}}(W)}{\operatorname{deg^{max}}(W)}\geq\rho$. Then for each $x\in X$ the probability that a particle type $x$ in the branching process $\mathfrak{X}_{W^{\dagger}}$ has no children is at least $g_{\rho}$.
Proof.
Indeed, the number of children of such a particle has distribution $\mathsf{Poi}(\int_{z\in X}\frac{W(x,z)}{\deg_{W}(z)})$. The lemma follows by noting that $\int_{z\in X}\frac{W(x,z)}{\deg_{W}(z)}\leq\frac{1}{\rho}$. ∎
The next lemma tells us that in each generation the expected number of particles is 1. It actually gives a more precise description of the distribution of the particles.
Lemma 4.6.
Suppose that $W$ is a nondegenerate $L^{1}$-kernel. Consider the branching process $\mathfrak{X}_{W^{\dagger}}$. For $k\in\mathbb{N}_{0}$ and $A\in\mathcal{B}$, let $Y_{k,A}$ be the number of particles in generation $k$ and of type $A$. Then we have $\mathbf{E}[Y_{k,A}]=\mu(A)$.
Proof.
We prove the claim by induction. The base case $k=0$ is clear since the root of $\mathfrak{X}_{W^{\dagger}}$ is chosen according to $\mu$. Let us move to the induction step $k\to k+1$. The number of particles in generation $k+1$ of type in $A$ which are born from a given particle in generation $k$ of type $x\in X$ has distribution $\mathsf{Poi}\left(\int_{z\in A}\frac{W(x,z)}{\deg_{W}(z)}\right)$. In particular, the expected number of particles in generation $k+1$ of type in $A$ which are born from a given particle in generation $k$ of type $x$ is $\int_{z\in A}\frac{W(x,z)}{\deg_{W}(z)}$. By the induction hypothesis, the expected number of particles in generation $k+1$ of type in $A$ is obtained by integrating the above uniformly over the choice of $x$,
$\mathbf{E}[Y_{k+1,A}]$ $=\int_{x\in X}\int_{z\in A}\frac{W(x,z)}{\deg_{W}(z)}\ \mathsf{d}\mu(z)=\int_{z\in A}\frac{1}{\deg_{W}(z)}\int_{x\in X}W(x,z)\ \mathsf{d}\mu(z)$
$=\int_{z\in A}\frac{1}{\deg_{W}(z)}\cdot\deg_{W}(z)\ \mathsf{d}\mu(z)=\mu(A)\;,$
as was needed. ∎
Lemma 4.7.
Suppose that $W$ is a bounded-degree kernel with positive minimum degree. Then the branching process $\mathfrak{X}_{W^{\dagger}}$ goes extinct almost surely.
Proof.
For $k,j\in\mathbb{N}_{0}$, let $q_{k,j}$ be the probability that in $\mathfrak{X}_{W^{\dagger}}$ there are exactly $j$ particles in the $k$-th generation. By the law of total probability, we have for each $k\in\mathbb{N}$ that
$\sum_{j=0}^{\infty}q_{k,j}=1\;.$ (24)
We have from Lemma 4.6 that for each $k\in\mathbb{N}$, $\sum_{j=0}^{\infty}jq_{k,j}=1$. In particular, for every $J\in\mathbb{N}$,
$\sum_{j=1}^{J}q_{k,j}>1-q_{k,0}-\tfrac{1}{J}\;.$ (25)
The sequence $(q_{k,0})_{k=0}^{\infty}$ is nondecreasing and hence has a limit $L=\lim_{k\to\infty}q_{k,0}$. The lemma amounts to proving that $L=1$. Assume for contradiction that $L<1$. Let
$J:=\left\lceil\frac{20}{1-L}\right\rceil\;.$ (26)
Let $g>0$ be from Lemma 4.5 for our kernel $W$. Define $g^{*}:=g^{J}$. Note that for each $j\leq J$, $g^{*}$ is a lower bound for going extinct in generation $k+1$ ($k$ arbitrary) given that in generation $k$ there were $j$ particles.
Let $K\in\mathbb{N}$ be such that for every $k\geq K$ we have that
$q_{k,0}\in\left[L-\frac{g^{*}(1-L)}{10},L\right]\;.$ (27)
Note that we have 0 particles in a generation $k+1$ if there were 0 particles in generation $k$ or the number of particles in generation $k$ was between $1$ and $J$ and none of these particles had any children (this is obviously not if and only if). By the way we set $g^{*}$ we have
[…]
Take $k\geq K$ and use (27) for $q_{k,0}$ and $q_{k+1,0}$ in (28).
$L\geq(1-\tfrac{1}{J})g^{*}+(1-g^{*})\left(L-\frac{g^{*}(1-L)}{10}\right)\;.$
Pedestrian manipulations (in which we divide both sides by $g^{*}>0$) give
[…]
a contradiction. ∎
4.3 Weak isomorphism of kernels
Last, we will use the notion of weak isomorphism of $L^{\infty}$-kernels. While the original definition goes via homomorphism densities, here we will rather use another property. As is usual in the theory of graph limits, given a kernel $Z:X^{2}\to\mathbb{R}$ and measure preserving map $\pi:X\to X$, we write $Z^{\pi}$ for the kernel defined by $Z^{\pi}(x,y)=Z(\pi(x),\pi(y))$.
Lemma 4.8 (Corollary 10.35(a) in [20]).
Suppose that $Z_{1}$ and $Z_{2}$ are two weakly isomorphic $L^{\infty}$-kernels. Then there exist measure preserving maps $\pi_{1},\pi_{2}:X\to X$ such that $Z_{1}^{\pi_{1}}(x,y)=Z_{2}^{\pi_{2}}(x,y)$ for almost every $(x,y)\in X^{2}$.
4.4 More on fractional isomorphism and sub-sigma-algebras
Let $(X,\mathcal{B})$ be a standard Borel space endowed with a Borel probability measure $\mu$. Let $W$ be an $L^{2}$-kernel on $X$. The corresponding Hilbert–Schmidt integral operator $T_{W}\colon L^{2}(X,\mu)\to L^{2}(X,\mu)$ is defined by
$T_{W}(f)(x)=\int_{X}W(x,y)f(y)\mathsf{d}\mu(y)\;.$ (29)
We say that $\mathcal{C}\subset\mathcal{B}$ is a $\mu$-relatively complete sub-sigma-algebra of $\mathcal{B}$ if it is a sub-sigma-algebra such that every $Z\in\mathcal{B}$ satisfies $Z\in\mathcal{C}$ whenever there is $Z_{0}\in\mathcal{C}$ such that $\mu(Z\triangle Z_{0})=0$. Let $\Theta_{\mu}$ be the collection of all $\mu$-relatively complete sub-sigma-algebras of $\mathcal{B}$. For $\mathcal{X}\subset\mathcal{B}$ we write $\langle\mathcal{X}\rangle$ for the unique relatively complete sub-sigma-algebra generated by $\mathcal{X}$; this notion is well-defined as a consequence of [11, Claim 5.4].
For an $L^{\infty}$-akernel $W$, we recall from [11, Section 5] the construction of a sequence $(\mathcal{C}^{W}_{n})_{n\in\mathbb{N}_{0}}$ of non-decreasing sigma-algebras which are all subalgebras of $\mathcal{B}$. We begin with some necessary definitions. Let $\mathcal{C}\in\Theta_{\mu}$. For $\mathcal{D}\in\Theta_{\mu}$ we say that $(\mathcal{C},\mathcal{D})$ is $W$-invariant if $T_{W}(L^{2}(X,\mathcal{C},\mu))\subset L^{2}(X,\mathcal{D},\mu)$; in particular, we say that $\mathcal{C}$ is $W$-invariant if $(\mathcal{C},\mathcal{C})$ is $W$-invariant. Let $Z_{\mathcal{C}}$ be the collection of $\mathcal{D}\in\Theta_{\mu}$ such that $(\mathcal{C},\mathcal{D})$ is $W$-invariant. Write $m(\mathcal{C})$ for the collection of sets $S\in\mathcal{B}$ such for all $\mathcal{D}\in Z_{\mathcal{C}}$ we have $S\in\mathcal{D}$. Now the collection $(\mathcal{C}^{W}_{n})_{n\in\mathbb{N}_{0}}$ of sigma-algebras is given follows. Let $\mathcal{C}^{W}_{0}=\langle\{\varnothing,X\}\rangle$ and inductively define $\mathcal{C}^{W}_{n}=m(\mathcal{C}^{W}_{n-1})$. Finally, define $\mathcal{C}(W)=\left\langle\bigcup_{n\in\mathbb{N}_{0}}\mathcal{C}^{W}_{n}\right\rangle$ as the unique relatively complete sigma-algebra generated by the union of these sigma-algebras.
The following lemma is a straight-forward generalization of [11, Corollary 6.7].
Lemma 4.9.
Suppose $W$ is an $L^{\infty}$-akernel with maximum degree $c$. Then $i_{W}$ is measurable and we have
$\langle\{i_{W}^{-1}(A)\mid A\in\mathcal{B}(\mathbb{P}^{c})\}\rangle=\mathcal{C}(W),$
i. e., the minimum relatively complete sub-sigma-algebra of $\mathcal{B}$ that makes the map $i_{W}$ measurable is $\mathcal{C}(W)$.
4.5 Conditional expectation
Let $R$ be an integrable random variable on a standard Borel space $(X,\mathcal{B})$ equipped with a Borel probability measure $\mu$. Recall that $\Theta_{\mu}$ is the collection of all $\mu$-relatively complete sub-sigma-algebras of $\mathcal{B}$. Let $\mathcal{C}\in\Theta_{\mu}$. The conditional expectation $\mathbf{E}[R|\mathcal{C}]$ of $R$ given $\mathcal{C}$ is a $\mathcal{C}$-measurable integrable random variable on $(X,\mathcal{B},\mu)$ such that for all $A\in\mathcal{C}$ we have
$\int_{A}\mathbf{E}[R|\mathcal{C}]\mathsf{d}\mu=\int_{A}R\mathsf{d}\mu\;.$ (30)
The following is a useful standard fact about conditional expectations.
Fact 4.10.
Let $R$ and $S$ be integrable random variables on a standard Borel space $(X,\mathcal{B})$ equipped with a Borel probability measure $\mu$. Let $\mathcal{C}\in\Theta_{\mu}$. Suppose that $RS$ is integrable and $R$ is $\mathcal{C}$-measurable. Then $\mathbf{E}[RS|\mathcal{C}]=R\mathbf{E}[S|\mathcal{C}]$ holds $\mu$-almost everywhere.
5 A factorization result for fractional isomorphism
This section deals with Theorem 5.1, which is a factorization result for fractional isomorphism of disconnected kernels. It says that two kernels $U$ and $W$, are fractionally isomorphic if and only if, for each connected kernel $\Gamma$, the total measure of connected components which are fractionally isomorphic to $\Gamma$ is the same in both $U$ and $W$. This result allows us to extend Theorem 1.7 to disconnected kernels in Theorem 6.1.
To state the result and also for subsequent parts of the paper, it is useful to introduce the concept of subgraphons or subkernels. Given a kernel $U$ on $X$, and $Y\subset X$ measurable with $\mu(Y)>0$, we define $U_{\restriction Y}$ as the restriction of $U$ to $Y\times Y$. We view $U_{\restriction Y}$ again as a kernel, that is, we equip $Y$ with the probability measure $\frac{\mu(Y\cap\cdot)}{\mu(Y)}$. Next, we introduce a rescaling $U\llbracket Y\rrbracket:Y\times Y\rightarrow\mathbb{R}$ defined by $U\llbracket Y\rrbracket(x,y)=\mu(Y)U_{\restriction Y}(x,y)$. This rescaling is chosen so that if $Y$ is a connected component of $U$ (or a union of connected components) then for every $y\in Y$ we have $\deg_{U\llbracket Y\rrbracket}(y)=\deg_{U}(y)$. More generally, for each $y\in Y$ we have $i_{U\llbracket Y\rrbracket}(y)=i_{U}(y)$ and for every $A\in\mathcal{B}(\mathbb{P}^{c})$ we have
$\mu(Y)\nu_{U\llbracket Y\rrbracket}(A)=\mu\left(i_{U}^{-1}(A)\cap Y\right)\;.$ (31)
The main result now reads as follows.
Theorem 5.1.
Suppose that $U$ and $W$ are two $L^{\infty}$-kernels. Let $X=\Lambda_{0}\sqcup\bigsqcup_{i=1}^{N_{U}}\Lambda_{i}$ and $X=\Omega_{0}\sqcup\bigsqcup_{i=1}^{{N_{W}}}\Omega_{i}$ be decompositions into connected components of $U$ and of $W$, respectively, with $\Lambda_{0}$ and $\Omega_{0}$ being the sets of isolated elements. Here, $N_{U},N_{W}\in\mathbb{N}\cup\{\infty\}$ and they are not necessarily equal. Then $U$ and $W$ are fractionally isomorphic if and only if for each connected kernel $\Gamma$ we have
$\sum_{i\geq 1:\text{$U\llbracket\Lambda_{i}\rrbracket$ is frac.\ iso.\ to $\Gamma$}}\mu(\Lambda_{i})=\sum_{i\geq 1:\text{$W\llbracket\Omega_{i}\rrbracket$ is frac.\ iso.\ to $\Gamma$}}\mu(\Omega_{i})\;.$ (32)
Note that if (32) holds, then $\mu(\Lambda_{0})=\mu(\Omega_{0})$. Indeed, there are at most countably many mutually fractionally nonisomorphic akernels $\Gamma$ for which the value in (32) is positive. The sum of these values is, when reasoning from the left-hand side, $1-\mu(\Lambda_{0})$, and when reasoning from the right-hand side, $1-\mu(\Omega_{0})$.
While the proof of Theorem 5.1 is not long, a reader without much experience with graphon fractional isomorphism may not find it illuminating. For propedeutic reasons, let us first prove an analogous theorem for finite graphs, a result which also appears new but we do not need it otherwise. In Section 1.1, we introduced the definition of fractional isomorphism of graphs based on iterated degree sequences. Another definition, based on [23], is more convenient here, and we state it with an additional twist. Suppose that $G$ is a graph and $\mathcal{P}=\{P_{j}\}_{j=1}^{k}$ is a partition of $V(G)$ into nonempty sets. Suppose that $\mathbf{D}=(D_{j,\ell})_{j,\ell=1}^{k}$ is a matrix of integers, and $\mathbf{p}=(p_{j})_{j=1}^{k}$ is a vector of positive reals. We say that $\mathcal{P}$ is an equitable partition for template $(\mathbf{D},\mathbf{p})$ if for every $j,\ell\in[k]$ we have
$\frac{|P_{j}|}{v(G)}$ $=p_{j}\quad\mbox{, and }$ (33)
$\deg_{G}(v,P_{\ell})$ $=D_{j,\ell}\quad\mbox{for every $v\in P_{j}$.}$ (34)
Furthermore, we say that $\mathcal{P}$ is the coarsest equitable partition of $G$ if every other equitable partition refines it. It is a well-known fact and also easy to check that every graph has a unique (up to the order of the cells) coarsest equitable partition.
We say that two graphs $G$ and $H$ are fractionally isomorphic if there exist $k\in\mathbb{N}$, and $\mathbf{D}$ and $\mathbf{p}$ as above so that both $G$ and $H$ have an equitable partition for template $(\mathbf{D},\mathbf{p})$. Again, it is well-known and easy to check that this is equivalent to the coarsest equitable partitions of $G$ and of $H$ having the same template.
The ‘additional twist’ in our definition is that (33) is phrased in terms of proportions rather than the usual cardinality condition $|P_{j}|=p_{j}$. So, while traditionally, two fractionally isomorphic graphs needed to have the same number of vertices, now, for example, each two three-regular graphs are fractionally isomorphic, even if their orders differ. [^5] Take $k=1$, $D_{1,1}=3$, $p_{1}=1$, and trivial equitable partitions $\{V(G)\}$, and $\{V(H)\}$.
We remark that fractional isomorphism of two finite graphs $G$ and $H$ is not equivalent to the fractional isomorphism of two step graphons $W_{G}$ and $W_{H}$ representing those graphs. For example, graphon representations of two three-regular graphs of different orders are not fractionally isomorphic.
However, by our choice of generalization, the factorization result for graphs has a particularly elegant form.
Theorem 5.2.
Suppose that $G$ and $H$ are two graphs. Let $V(G)=A_{1}\sqcup A_{2}\sqcup\ldots\sqcup A_{n}$ and $V(H)=B_{1}\sqcup B_{2}\sqcup\ldots\sqcup B_{m}$ be the connected components of $G$ and of $H$, respectively. Then $G$ and $H$ are fractionally isomorphic if and only if for each connected graph $\Gamma$ we have
$\frac{1}{v(G)}\cdot\sum_{i:\text{$G[A_{i}]$ is frac.\ iso.\ to $\Gamma$}}|A_{i}|=\frac{1}{v(H)}\cdot\sum_{i:\text{$H[B_{i}]$ is frac.\ iso.\ to $\Gamma$}}|B_{i}|\;.$ (35)
Proof.
We start with the routine ()($\Leftarrow$)( ⇐ ) direction. The idea is to stitch together corresponding cells of equitable partitions of components of $G[A_{i}]$ (and later $H[B_{i}]$) that are fractionally isomorphic to the same connected graph $\Gamma$. Let $\mathcal{R}$ be a family of arbitrary representatives of fractional isomorphism classes of those connected graphs which appear among components of $G$ (and equivalently, by 35, among components of $H$). That is, for each $G[A_{i}]$ there is a unique $\Gamma\in\mathcal{R}$ such that $G[A_{i}]$ and $\Gamma$ are fractionally isomorphic. For each $\Gamma\in\mathcal{R}$, let $(\mathbf{D}^{\Gamma},\mathbf{p}^{\Gamma})$ be the template of the coarsest equitable partition of $\Gamma$. We write $\mathbf{D}^{\Gamma}=(D^{\Gamma}_{j,\ell})_{j,\ell=1}^{k_{\Gamma}}$ and $\mathbf{p}^{\Gamma}=(p^{\Gamma}_{j})_{j=1}^{k_{\Gamma}}$. Let us now define a template $(\mathbf{D},\mathbf{p})$ of dimension $k=\sum_{\Gamma\in\mathcal{R}}k_{\Gamma}$. Rather than indexing the entries of the template by elements of $\{1,\ldots,k\}$, it is convenient to index them by $S=\{(\Gamma,j):\Gamma\in\mathcal{R},j\in[k_{\Gamma}]\}$. That is, we define $\mathbf{D}=(D_{s,q})_{s,q\in S}$ and $\mathbf{p}=(p_{s})_{s\in S}$ by
$D_{(\Gamma_{1},j_{1}),(\Gamma_{2},j_{2})}$ $=D^{\Gamma_{1}}_{j_{1},j_{2}}$ if $\text{if $\Gamma_{1}=\Gamma_{2}$},$
$D_{(\Gamma_{1},j_{1}),(\Gamma_{2},j_{2})}$ $=0$ if $\text{if $\Gamma_{1}\neq\Gamma_{2}$},$
$p_{(\Gamma,j)}$ $=\tfrac{\sum_{i:\text{$G[A_{i}]$ is frac.\ iso.\ to $\Gamma$}}|A_{i}|}{v(G)}\cdot p^{\Gamma}_{j}\;.$ (36)
We now define a partition $\mathcal{P}=(P_{s})_{s\in S}$ of $V(G)$ as follows,
$P_{(\Gamma,j)}=\bigcup_{i:\text{$G[A_{i}]$ is frac.\ iso.\ to $\Gamma$}}P^{i}_{j}\;,$
where $\mathcal{P}^{i}=(P^{i}_{j})_{j=1}^{k_{\Gamma}}$ is a partition of $A_{i}$ corresponding to template $(\mathbf{D}^{\Gamma},\mathbf{p}^{\Gamma})$, which exists as $\Gamma$ and $G[A_{i}]$ are fractionally isomorphic. It is routine to check that $\mathcal{P}$ is an equitable partition for template $(\mathbf{D},\mathbf{p})$. We can now repeat the whole construction for the graph $H$. The template we create this way is the same; this crucial fact uses (35) in (36). We conclude that $G$ and $H$ are fractionally isomorphic.
The bulk of the work is the ()($\Rightarrow$)( ⇒ ) direction. Assume that $G$ and $H$ are fractionally isomorphic. Let $(\mathbf{D},\mathbf{p})$ be their common template, and let $\mathcal{P}=\{P_{j}\}_{j=1}^{k}$ and $\mathcal{Q}=\{Q_{j}\}_{j=1}^{k}$ be the corresponding equitable partitions for $G$ and for $H$, respectively.
Consider an auxiliary graph $\Delta$ on vertex set $[k]$. Make a pair $j\ell$ an edge of $\Delta$ if $D_{j,\ell}>0$ (observe that this is equivalent to $D_{\ell,j}>0$). For a set $S\subset[k]$ write $V^{(G)}_{S}=\bigcup_{j\in S}P_{j}$ and $V^{(H)}_{S}=\bigcup_{j\in S}Q_{j}$.
Claim A.
For each connected component $A_{i}$ of $G$ there exists a connected component $C$ of $\Delta$ such $A_{i}\subset V^{(G)}_{C}$. Furthermore, for each $j,\ell\in C$, we have $\frac{|A_{i}\cap P_{j}|}{|P_{j}|}=\frac{|A_{i}\cap P_{\ell}|}{|P_{\ell}|}$.
Let $C$ be a connected component of $\Delta$ for which $V^{(G)}_{C}$ contains at least 1 vertex of $A_{i}$. To prove the first part of the statement, we need to prove that for any other connected component $C^{\prime}$ of $\Delta$, we have $e_{G}(A_{i}\cap V^{(G)}_{C},A_{i}\cap V^{(G)}_{C^{\prime}})=0$. Indeed, we have
$e_{G}(A_{i}\cap V^{(G)}_{C},A_{i}\cap V^{(G)}_{C^{\prime}})\leq e_{G}(V^{(G)}_{C},V^{(G)}_{C^{\prime}})=\sum_{j\in C}\sum_{\ell\in C^{\prime}}e_{G}(P_{j},P_{\ell})=0\;,$
where the last equality uses (34) and the fact that all the corresponding numbers $D_{j,\ell}$ are 0.
For the furthermore part, observe that it is enough to prove the equality $\frac{|A_{i}\cap P_{j}|}{|P_{j}|}=\frac{|A_{i}\cap P_{\ell}|}{|P_{\ell}|}$ only for those $j,\ell\in C$ that form an edge in $\Delta$. Using the fact that there are no edge between $A_{i}\cap P_{j}$ and $P_{\ell}\setminus A_{i}$ we have
[…]
and similarly
$e_{G}(A_{i}\cap P_{j},A_{i}\cap P_{\ell})=D_{\ell,j}|A_{i}\cap P_{\ell}|\;.$ (38)
We also have
$D_{j,\ell}|P_{j}|=e_{G}(P_{j},P_{\ell})=D_{\ell,j}|P_{\ell}|\;.$ (39)
Putting equations (37), (38) and (39) together (using that we may divide by $e_{G}(P_{j},P_{\ell})\neq 0$), we get the desired statement.
Claim B.
Suppose that $A_{i}$ is a connected component of $G$ and $C$ is a connected component in $\Delta$ such $A_{i}\subset V^{(G)}_{C}$. Then $G[A_{i}]$ is fractionally isomorphic to $G[V^{(G)}_{C}]$.
Proof.
We show that $\{A_{i}\cap P_{j}\}_{j\in C}$ and $\{P_{j}\}_{j\in C}$ are equitable partitions for $G[A_{i}]$ and for $G[V^{(G)}_{C}]$ with the same template. For (33), we proceed as follows. Write $\xi$ for the positive real number equal to $\frac{|A_{i}\cap P_{t}|}{|P_{t}|}$ for all $t\in C$; this is well-defined by Claim A. Now for each $j\in C$ we have
$\frac{|A_{i}\cap P_{j}|}{|A_{i}|}=\frac{|A_{i}\cap P_{j}|}{\sum_{t\in C}|A_{i}\cap P_{t}|}=\frac{\xi|P_{j}|}{\sum_{t\in C}\xi|P_{t}|}=\frac{|P_{j}|}{|V^{(G)}_{C}|}\;.$
This shows that the corresponding ratios in (33) are equal for the partition $\{A_{i}\cap P_{j}\}_{j\in C}$ in the graph $G[A_{i}]$ and for the partition $\{P_{j}\}_{j\in C}$ in the graph $G[V^{(G)}_{C}]$. Let us turn to verifying (34). Let $j,\ell\in C$ be given. Let us start with the graph $G[A_{i}]$. Let $v\in A_{i}\cap P_{j}$. We have $\deg_{G[A_{i}]}(v,A_{i}\cap P_{\ell})=\deg_{G}(v,P_{\ell})=D_{j,\ell}$, where the first equality uses that $A_{i}$ is a connected component. The calculation for the graph $G[V^{(G)}_{C}]$ is even more straightforward. Let $v\in P_{j}$. We have $\deg_{G[V^{(G)}_{C}]}(v,P_{\ell})=\deg_{G}(v,P_{\ell})=D_{j,\ell}$, as $V^{(G)}_{C}$ is a union of connected components of $G$ that includes $v$. ∎
We can now quickly finish the proof. Indeed, Claims A and B tell us that the left-hand side of (35) is equal to $\frac{1}{v(G)}\cdot\sum_{C}|V^{(G)}_{C}|$, where the summation runs over connected components $C$ of $\Delta$ such that $G[V^{(G)}_{C}]$ is fractionally isomorphic to $\Gamma$. By repeating the argument for $H$, we see that the right-hand side of (35) is equal to $\frac{1}{v(H)}\cdot\sum_{C}|V^{(H)}_{C}|$, where the summation runs over connected components $C$ of $\Delta$ such that $H[V^{(H)}_{C}]$ is fractionally isomorphic to $\Gamma$. The last observation we need to equate both sums, which follows easily from the common template, is that for any nonempty set $S\subset[k]$, the graphs $G[V^{(G)}_{S}]$ and $H[V^{(H)}_{S}]$ are fractionally isomorphic and $\frac{|V^{(G)}_{S}|}{v(G)}=\frac{|V^{(H)}_{S}|}{v(H)}$. ∎
In kernels, the role of the equitable partition is replaced by the sigma-algebra $\mathcal{C}(W)$ which we reintroduced in Section 4.4. The proof of Theorem 5.1 proceeds by adopting this formalism, but main ideas remain.
Proof of Theorem 5.1.
We start with the routine ()($\Leftarrow$)( ⇐ ) direction. The measure $\nu_{U}$ can be expressed as a convex combination of measures $\{\nu_{U\llbracket\Lambda_{i}\rrbracket}\}_{i=1}^{N_{U}}$,
$\nu_{U}=\sum_{i=1}^{N_{U}}\mu(\Lambda_{i})\cdot\nu_{U\llbracket\Lambda_{i}\rrbracket}+\mu(\Lambda_{0})\cdot\mathsf{Dirac}(\oslash)\;,$ (40)
where $\oslash\in\mathbb{P}^{0}$ is the trivial measure. In analogy with the proof of Theorem 5.2, let $\mathcal{R}$ be an arbitrary family of representatives of fractional isomorphism classes of connected kernels appearing among components of $U$. If $U\llbracket\Lambda_{i}\rrbracket$ is fractionally isomorphic to $\Gamma\in\mathcal{R}$ then $\nu_{U\llbracket\Lambda_{i}\rrbracket}=\nu_{\Gamma}$. Hence, we can rewrite (40),
$\nu_{U}=\sum_{\Gamma\in\mathcal{R}}\left(\sum_{i\geq 1:\text{$U\llbracket\Lambda_{i}\rrbracket$ is frac.\ iso.\ to $\Gamma$}}\mu(\Lambda_{i})\right)\cdot\nu_{\Gamma}+\mu(\Lambda_{0})\cdot\mathsf{Dirac}(\oslash)\;.$
We can now repeat the whole process for the kernel $W$ and end up with
$\nu_{W}=\sum_{\Gamma\in\mathcal{R}}\left(\sum_{i\geq 1:\text{$W\llbracket\Lambda_{i}\rrbracket$ is frac.\ iso.\ to $\Gamma$}}\mu(\Omega_{i})\right)\cdot\nu_{\Gamma}+\mu(\Omega_{0})\cdot\mathsf{Dirac}(\oslash)\;.$
We use (32) to conclude that $\nu_{U}=\nu_{W}$. Hence, $U$ and $W$ are fractionally isomorphic.
Let us now turn to the ()($\Rightarrow$)( ⇒ ) direction. Write $\Delta:=\mathbf{E}[U|\mathcal{C}(U)\times\mathcal{C}(U)]$. By Claim 5.7 in [11] we have
$\Delta=\mathbf{E}[U|\mathcal{B}(X)\times\mathcal{C}(U)]=\mathbf{E}[U|\mathcal{C}(U)\times\mathcal{B}(X)]\;.$ (41)
Claim A.
For each connected component $\Lambda_{i}$ of $U$ there exists a connected component $C$ of $\Delta$ such $\Lambda_{i}\subset C$ (up to a $\mu$-nullset). Furthermore, let $h_{i}:=\mathbf{E}[\mathbbm{1}_{\Lambda_{i}}|\mathcal{C}(U)]$. Then we have that $h_{i}=\frac{\mu(\Lambda_{i})}{\mu(C)}\cdot\mathbbm{1}_{C}$.
Proof.
Let $C$ be a connected component of $\Delta$ such that $\mu(\Lambda_{i}\cap C)>0$ (such a connected component exists). In order to prove the first part of the statement, we need to prove that for any other connected component $C^{\prime}$ of $\Delta$, we have $\int_{(\Lambda_{i}\cap C)\times(\Lambda_{i}\cap C^{\prime})}U(x,y)\mathsf{d}\mu^{2}(x,y)=0$. Indeed, we have
$0\leq\int_{(\Lambda_{i}\cap C)\times(\Lambda_{i}\cap C^{\prime})}U(x,y)\mathsf{d}\mu^{2}(x,y)\leq\int_{C\times C^{\prime}}U(x,y)\mathsf{d}\mu^{2}(x,y)=\int_{C\times C^{\prime}}\Delta(x,y)\mathsf{d}\mu^{2}(x,y)=0\;.$
Let us turn to the furthermore part. We establish the claim in three parts:
  1. (p1)
    $h_{i}$ is constant-0 almost everywhere outside $C$,
  2. (p2)
    $h_{i}$ is constant almost everywhere on $C$, and
  3. (p3)
    the value of $h_{i}$ on $C$ is $\frac{\mu(\Lambda_{i})}{\mu(C)}$.
Proof of (p1). This is obvious as $\Lambda_{i}\subset C$.
Proof of (p2). Suppose for contradiction that $h_{i}$ is not constant on $C$. That means that there exists $r\geq 0$ such that the sets $R^{+}=\{x\in C\mid h_{i}(x)>r\}$ and $R^{-}=\{x\in C\mid h_{i}(x)\leq r\}$ both have positive measure. As $h_{i}$ is the conditional expectation with respect to $\mathcal{C}(U)$, it is $\mathcal{C}(U)$-measurable. In particular, $R^{+}$ and $R^{-}$ are $\mathcal{C}(U)$-measurable. Note also that since $R^{+}\sqcup R^{-}$ is a nontrivial partition of a connected component of $\Delta$, we have
$\int_{R^{+}\times R^{-}}\Delta(x,y)\mathsf{d}\mu^{2}(x,y)>0\;.$ (42)
Using that $\Lambda_{i}$ is a connected component of $U$, we have
$\begin{split}\int_{R^{+}\times R^{-}}\mathbbm{1}_{\Lambda_{i}}(x)U(x,y)\mathsf{d}\mu^{2}(x,y)&=\int_{R^{+}\times R^{-}}\mathbbm{1}_{\Lambda_{i}}(x)\mathbbm{1}_{\Lambda_{i}}(y)U(x,y)\mathsf{d}\mu^{2}(x,y)\\ &=\int_{R^{+}\times R^{-}}\mathbbm{1}_{\Lambda_{i}}(y)U(x,y)\mathsf{d}\mu^{2}(x,y)\;.\end{split}$ (43)
Let us express the first integral of (43).
$\int_{R^{+}\times R^{-}}\mathbbm{1}_{\Lambda_{i}}(x)U(x,y)\mathsf{d}\mu^{2}(x,y)$ $=\int_{R^{+}}\mathbbm{1}_{\Lambda_{i}}(x)\int_{R^{-}}U(x,y)\mathsf{d}\mu(y)\mathsf{d}\mu(x)$
(41), and $R^{-}\in\mathcal{C}(U)$ $=\int_{R^{+}}\mathbbm{1}_{\Lambda_{i}}(x)\int_{R^{-}}\Delta(x,y)\mathsf{d}\mu(y)\mathsf{d}\mu(x)\;.$
Since the function $x\mapsto\int_{R^{-}}\Delta(x,y)\mathsf{d}\mu(y)$ is $\mathcal{C}(U)$-measurable by (41), by Fact 4.10 we have
$\int_{R^{+}}\mathbbm{1}_{\Lambda_{i}}(x)\int_{R^{-}}\Delta(x,y)\mathsf{d}\mu(y)\mathsf{d}\mu(x)=\int_{R^{+}}h_{i}(x)\int_{R^{-}}\Delta(x,y)\mathsf{d}\mu(y)\mathsf{d}\mu(x)\;.$
That is, we have derived $\int_{R^{+}\times R^{-}}\mathbbm{1}_{\Lambda_{i}}(x)U(x,y)\mathsf{d}\mu^{2}(x,y)=\int_{R^{+}\times R^{-}}h_{i}(x)\Delta(x,y)\mathsf{d}\mu^{2}(x,y)$. Similarly, starting with the last integral of (43), we get $\int_{R^{+}\times R^{-}}\mathbbm{1}_{\Lambda_{i}}(y)U(x,y)\mathsf{d}\mu^{2}(x,y)=\int_{R^{+}\times R^{-}}h_{i}(y)\Delta(x,y)\mathsf{d}\mu^{2}(x,y)$. Substituting these to (43), we get
$\int_{R^{+}\times R^{-}}h_{i}(x)\Delta(x,y)\mathsf{d}\mu^{2}(x,y)=\int_{R^{+}\times R^{-}}h_{i}(y)\Delta(x,y)\mathsf{d}\mu^{2}(x,y)\;.$
We have $h_{i}(x)>r$, $h_{i}(y)\leq r$, and further these terms $h_{i}(x)$ and $h_{i}(y)$ multiply a nonzero main term by (42). This is a contradiction.
Proof of (p3). Let $H$ be the value of $h_{i}$ on $C$. The definition of $h_{i}$ via the conditional expectation gives $\int_{X}h_{i}\mathsf{d}\mu=\int_{X}\mathbbm{1}_{\Lambda_{i}}\mathsf{d}\mu$. Together with $\int_{X}h_{i}\mathsf{d}\mu=H\mu(C)$ and $\int_{X}\mathbbm{1}_{\Lambda_{i}}\mathsf{d}\mu=\mu(\Lambda_{i})$, the claim follows. ∎
Claim B.
For a connected component $\Lambda_{i}$ of $U$ and the connected component $C$ of $\Delta$ such $\Lambda_{i}\subset C$ (up to a $\mu$-nullset), we have that $U\llbracket\Lambda_{i}\rrbracket$ is fractionally isomorphic to $\Delta\llbracket C\rrbracket$.
Proof.
We verify fractional isomorphism using Definition 2.3. Let $A\subset\mathbb{P}^{c}$ be arbitrary. By (31) we have
$\mu(\Lambda_{i})\nu_{U\llbracket\Lambda_{i}\rrbracket}(A)=\mu\left(i_{U}^{-1}(A)\cap\Lambda_{i}\right)=\int_{i_{U}^{-1}(A)}\mathbbm{1}_{\Lambda_{i}}(x)\mathsf{d}\mu(x)\;.$
The set $i_{U}^{-1}(A)$ is $\mathcal{C}(U)$-measurable by Lemma 4.9. Hence we have
[…]
as desired. ∎
We can now quickly finish the proof. Indeed, Claims A and B tell us that for the left-hand side of (32) we have
[…]
where the summation runs over connected components $C$ of $\Delta$ such that $\Delta\llbracket C\rrbracket$ is fractionally isomorphic to $\Gamma$. By repeating the argument for $W$ and for $\Delta^{\prime}:=\mathbf{E}[W|\mathcal{C}(W)\times\mathcal{C}(W)]$, we see that
[…]
where the summation runs over connected components $C$ of $\Delta^{\prime}$ such that $\Delta^{\prime}\llbracket C\rrbracket$ is fractionally isomorphic to $\Gamma$. We now use one of the main characterizations of fractional isomorphism. Namely, Theorem 1.2(3) in [11] tells us that kernels $U/{\mathcal{C}(U)}$ and $W/{\mathcal{C}(W)}$ (introduced therein) are isomorphic. By routinely expanding the definitions of $U/{\mathcal{C}(U)}$ and $W/{\mathcal{C}(W)}$, this implies that $\Delta$ and $\Delta^{\prime}$ are weakly isomorphic. By Lemma 4.8, there exist measure preserving maps $\pi$ and $\pi^{\prime}$ such that $\Delta^{\pi}=(\Delta^{\prime})^{\pi^{\prime}}$ almost everywhere. We replace the host kernel $\Delta$ by $\Delta^{\pi}$ in (44). Obviously, the sum, which we call $\sum_{C:(*mpm)}\mu(C)$ has not changed. Similarly, we write $\sum_{C:(**mpm)}\mu(C)$ has not for (45) in which $\Delta^{\prime}$ is replaced by $(\Delta^{\prime})^{\pi^{\prime}}$. We have
[…]
as was needed. ∎
6 Theorem 1.7: generalization and proof
6.1 Generalization of Theorem 1.7 to disconnected kernels
Theorem 1.7 does not generalize verbatim to disconnected kernels. To see this, partition $X=X_{1}\sqcup X_{2}$ with $\mu(X_{1})=0.2$ and $\mu(X_{2})=0.8$. Let $W$ be constant-13 on $X_{1}\times X_{1}$, constant-7 on $X_{2}\times X_{2}$, and constant-0 otherwise. No matter whether $x\in X$ from Definition 1.5 is sampled from $X_{1}$ or from $X_{2}$, the entire branching process $\mathfrak{U}_{W}$ will stay confined to that connected component in which it will behave as in $U\equiv 1$. Below, we introduce a notion of piecewise projective fractional isomorphism which exactly captures the fact that the projective constants may be different on different connected components.
The notion of projective fractional isomorphism extends to partial kernels. More precisely, take sets $X_{A},X_{B}\subset X$ of positive measure and two symmetric measurable functions $A:X_{A}^{2}\rightarrow[0,\infty)$ and $B:X_{B}^{2}\rightarrow[0,\infty)$ with $\operatorname{ess\,sup}_{x\in X_{A}}\int_{y\in X_{A}}A(x,y),\operatorname{ess\,sup}_{x\in X_{B}}\int_{y\in X_{B}}B(x,y)<\infty$. We say that $A$ and $B$ are projectively fractionally isomorphic if they are projectively fractionally isomorphic when viewed as kernels, that is, $A$ as a function on the square of the set $X_{A}$ equipped with the probability measure $\frac{\mu_{\restriction{X_{A}}}(\cdot)}{\mu(X_{A})}$, and $B$ as a function on the square of the set $X_{B}$ equipped with the probability measure $\frac{\mu_{\restriction{X_{B}}}(\cdot)}{\mu(X_{B})}$. Note that the fact that the ground sets are different is not an issue for the definition as fractional isomorphism is defined using measures in $\mathcal{M}_{=1}(\mathbb{P}^{c})$, a space independent of the ground set.
Suppose that $U$ and $W$ are nondegenerate bounded-degree kernels. Let us consider a decomposition $X=\Lambda_{1}\sqcup\Lambda_{2}\sqcup\ldots$ into connected components of $U$, and a decomposition $X=\Omega_{1}\sqcup\Omega_{2}\sqcup\ldots$ into connected components of $W$. We say that $U$ and $W$ are piecewice projectively fractionally isomorphic if for each connected bounded-degree kernel $\Gamma$ we have
$\sum_{i:\text{$U_{\restriction\Lambda_{i}}$ is proj.frac.\ iso.\ to $\Gamma$}}\mu(\Lambda_{i})=\sum_{i:\text{$W_{\restriction\Omega_{i}}$ is proj.frac.\ iso.\ to $\Gamma$}}\mu(\Omega_{i})\;.$ (46)
Obviously, piecewice projective fractional isomorphism is an equivalence relation on nondegenerate bounded-degree kernels. We can now formulate the generalization of Theorem 1.7 to disconnected kernels.
Theorem 6.1.
Suppose that $U$ and $W$ are $L^{\infty}$-kernels with positive minimum degrees. Then $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have the same distribution if and only if $U$ and $W$ are piecewise projectively fractionally isomorphic.
Proof.
The ()($\Leftarrow$)( ⇐ ) direction, while somewhat tedious, follows by expanding the definitions. We postpone this verification to Section 6.2. Here, we focus on the ()($\Rightarrow$)( ⇒ ) direction.
So, suppose that $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have the same distribution. We now claim that $\mathfrak{U}^{-}_{U}$ and $\mathfrak{U}^{-}_{W}$ have the same distributions. To illustrate a difficulty and a way to circumvent it, let us consider a toy example from the realm of real numbers rather than branching processes. Suppose that $a,b,a^{+},a^{-},b^{+},b^{-}\in\mathbb{R}$ are such that $a=a^{+}+a^{-}$ and $b=b^{+}+b^{-}$. Simply being told that $a=b$ alone is not adequate information to conclude that $a^{-}=b^{-}$. However, we do get this conclusion if we are also told that $a^{+},b^{+}\in\mathbb{Z}$ and $a^{-},b^{-}\in[0,1)$. We can apply a similar idea of splitting into big and small to the branching processes $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ whose equality in distribution we are given. Here, the ‘big parts’ are the ancestral parts and the ‘small parts’ are $\mathfrak{U}^{-}_{U}$ and $\mathfrak{U}^{-}_{W}$. The ancestral parts are infinite since they contain an infinite $\mathsf{anc}$-path. To complement this, Lemma 4.7 tells us that $\mathfrak{U}^{-}_{U}$ and $\mathfrak{U}^{-}_{W}$ are finite almost surely (here, we also use Fact 4.4 to transfer from $\mathfrak{X}_{U^{\dagger}}$ and $\mathfrak{X}_{W^{\dagger}}$ to $\mathfrak{U}^{-}_{U}$ and $\mathfrak{U}^{-}_{W}$). It then indeed follows that $\mathfrak{U}^{-}_{U}$ and $\mathfrak{U}^{-}_{W}$ have the same distribution.
Now, Theorem 1.3 applies and tells us that $U^{\dagger}$ and $W^{\dagger}$ are fractionally isomorphic. The last piece in the puzzle is the following crucial lemma.
Lemma 6.2.
Suppose that $U$ and $W$ are $L^{\infty}$-kernels with positive minimum degrees such that $U^{\dagger}$ and $W^{\dagger}$ are fractionally isomorphic. Then $U$ and $W$ are piecewise projectively fractionally isomorphic.
We give a proof of Lemma 6.2 in Section 6.3. ∎
6.2 Proof of the ()($\Leftarrow$)( ⇐ ) direction of Theorem 6.1
Given any bounded-degree kernel $W$ with positive minimum degree, we first give a construction of $i_{W^{\dagger}}$ from $i_{W}$. Then, in Lemmas 6.5 and 6.6, we conclude that $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have indeed the same distribution.
Suppose that $0<a\leq b<\infty$ are given. We will come up with a space, denoted $\mathbb{P}^{a,b}$, which is a subset of $\mathbb{P}^{b}$. Define $\mathbb{D}^{(0)}_{a,b}:=\{\alpha\in\mathbb{P}^{b}:D(\alpha)<a\}$, and inductively, $\mathbb{D}^{(\ell)}_{a,b}:=\left\{\alpha\in\mathbb{P}^{b}\mid\mu_{\alpha}(\mathbb{D}^{(\ell-1)}_{a,b})>0\right\}$. Let $\mathbb{P}^{a,b}:=\mathbb{P}^{b}\setminus\bigcup_{\ell=0}^{\infty}\mathbb{D}^{(\ell)}_{a,b}$. So, $\mathbb{P}^{a,b}$ are elements of $\mathbb{P}^{b}$ whose degree is at least $a$, and which only see elements of degrees at least $a$, and which only see elements which only see elements of degrees at least $a$, …. Obviously, this property holds for elements of the support of $\nu_{Z}$ of any akernel $Z$ with $\operatorname{deg^{max}}(Z)\leq b$ and $\operatorname{deg^{min}}(Z)\geq a$.
We now introduce a notion of degree tilting. For all $\alpha\in\mathbb{P}^{a,b}$ we define degree tilting $\alpha^{\ddagger}\in\mathbb{P}^{b/a,a/b}$ by $\alpha^{\ddagger}(0)=\star$ and inductively for $n\in\mathbb{N}$ defined for each Borel $A\subset\mathbb{P}_{n-1}^{b/a,a/b}$ by
$\alpha^{\ddagger}(n+1)(A)=\int_{\{\gamma\in\mathbb{P}^{a,b}\mid\gamma^{\ddagger}(n-1)\in A\}}\frac{1}{D(\beta)}\mathsf{d}\mu_{\alpha}(\beta)\;.$
Note that this definition is not circular as we only need the $(n-1)$-st component of all $\gamma^{\ddagger}$ to define the $n$-th component of $\alpha^{\ddagger}$. It is a straightforward exercise, that indeed $\alpha^{\ddagger}\in\mathbb{P}^{b/a,a/b}$.
We now return to the kernel $W$. For each $x\in X$, define $\lambda_{x,W}\in\mathcal{M}_{\leq\Delta}(X)$ by $\lambda_{x,W}(Y)=\int_{Y}W(x,z)\mathsf{d}\mu(z)$.
Lemma 6.3.
For each $x\in X$, $\mu_{i_{W}(x)}$ is the pushforward of $\lambda_{x,W}$ via $i_{W}$.
Proof.
As $\mathcal{B}(\mathbb{P}^{c})$ is generated by $\bigcup_{n\in\mathbb{N}}\{p_{n,\infty}^{-1}(A)\mid A\in\mathcal{B}(\mathbb{P}_{n}^{c})\}$, it suffices to show that $\mu_{i_{W}(x)}(p_{n,\infty}^{-1}(A))=(i_{W})_{*}\lambda_{W,x}(p_{n,\infty}^{-1}(A))$ for each $n\in\mathbb{N}$ and $A\in\mathcal{B}(\mathbb{P}_{n}^{c})$. We have
$\mu_{i_{W}(x)}(p_{n,\infty}^{-1}(A))$ $=(p_{n,\infty})_{*}\mu_{i_{W}(x)}(A)$
$=i_{W}(x)(n+1)(A)$
=$=\int_{\{y\in X\mid p_{n,\infty}(i_{W}(y))\in A\}}W(x,z)\mathsf{d}\mu(z)$
$=\int_{i_{W}^{-1}(p_{n,\infty}^{-1}(A))}W(x,z)\mathsf{d}\mu(z)$
$=\lambda_{W,x}(i_{W}^{-1}(p_{n,\infty}^{-1}(A)))$
$=((i_{W})_{*}\lambda_{W,x})(p_{n,\infty}^{-1}(A))\;,$
as was needed. ∎
Lemma 6.4.
For each $x\in X$ we have $i_{W^{\dagger}}(x)=(i_{W}(x))^{\ddagger}$.
Proof.
We proceed by induction on $n\in\mathbb{N}$. For $n=0$ we have $i_{W^{\dagger}}(0)=\star=(i_{W})^{\ddagger}(0)$. As an induction hypothesis assume that $i_{W^{\dagger}}(y)(k)=(i_{W}(y))^{\ddagger}(k)$ for each $y\in X$ and $k\in[n]_{0}$.
For each $A\in\mathcal{B}(\mathbb{P}_{n}^{\Delta/d})$, let $B_{n}:=\{\alpha\in\mathbb{P}^{\Delta,d}\mid\alpha^{\ddagger}(n)\in A\}$. For each $x\in X$,
$(i_{W^{\dagger}}(x)(n+1))(A)=$$\int_{\{y\in X\mid i_{W^{\dagger}}(y)(n)\in A\}}W^{\dagger}(x,z)\mathsf{d}\mu(z)$
[…]
$=$ $\int_{\{\alpha\in\mathbb{P}^{\Delta,d}\mid\alpha^{\ddagger}(n)\in A\}}\frac{1}{D(\beta)}\mathsf{d}\mu_{i_{W}(x)}(\beta)\;.$
Given an akernel $W$ with minimum degree $d>0$ and maximum degree $\Delta$, we introduce a stochastic process $\mathfrak{A}_{W}=(\mathfrak{A}^{(0)}_{W},\mathfrak{A}^{(1)}_{W},\ldots)\in(\mathbb{P}^{\Delta/d})^{\mathbb{N}_{0}}$ as follows. First, generate $z_{0}$ according to $\mu$. Given $z_{k-1}$, generate $z_{k}$ according to the intensity measure $\frac{W(z_{k-1},\cdot)}{\deg_{W}(z_{k-1})}$. Now, for each $k\in\mathbb{N}_{0}$, let $\mathfrak{A}^{(k)}_{W}:=\big{(}i_{W}(z_{k})\big{)}^{\ddagger}$.
The ()($\Leftarrow$)( ⇐ ) direction of Theorem 6.1 follows from the following two lemmas.
Lemma 6.5.
Suppose that $U$ and $W$ are bounded-degree kernels with positive minimum degrees. If $U$ and $W$ are piecewise projectively fractionally isomorphic then $\mathfrak{A}_{U}$ and $\mathfrak{A}_{W}$ have the same distribution.
Lemma 6.6.
Suppose that $U$ and $W$ are bounded-degree kernels with positive minimum degrees with the property that $\mathfrak{A}_{U}$ and $\mathfrak{A}_{W}$ have the same distribution. Then $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have the same distribution.
Proof of Lemma 6.5.
Let $(\Gamma_{j})_{j\in J}$ be a collection of connected mutually non-projectively fractionally isomorphic kernels for which (46) is positive, and let $f_{j}$ be the value of (46) for $\Gamma_{j}$. We have $\sum_{j\in J}f_{j}=1$. We couple the random choice of the root $z_{0}$ in $U$ and of the root $z^{\prime}_{0}$ in $W$ as follows. First, we pick a random $j\in J$ from the distribution $(f_{i})_{j\in J}$. Then, in $U$ we pick a random connected component $U_{\restriction\Lambda_{i}}$ that is projectively fractionally isomorphic to $\Gamma_{j}$. That is, $i$ is taken from the distribution
$\left(\frac{\mu(\Lambda_{i})}{f_{j}}\right)_{i:\text{$U_{\restriction\Lambda_{i}}$ is proj.frac.\ iso.\ to $\Gamma_{j}$}}\;.$
Similarly, in $W$ we pick a random connected component $W_{\restriction\Omega_{i^{\prime}}}$ that is projectively fractionally isomorphic to $\Gamma_{j}$. Write $U^{*}:=U_{\restriction\Lambda_{i}}/\|U_{\restriction\Lambda_{i}}\|_{1}$, $W^{*}:=W_{\restriction\Omega_{i^{\prime}}}/\|W_{\restriction\Omega_{i^{\prime}}}\|_{1}$. Last, we take $z_{0}\in\Lambda_{i}$ and $z^{\prime}_{0}\in\Omega_{i^{\prime}}$ uniformly at random but coupled in way so that
$i_{U^{*}}(z_{0})=i_{W^{*}}(z^{\prime}_{0})\;.$
This is possible because of projective fractional isomorphism. Note that $z_{0}$ and $z^{\prime}_{0}$ are uniformly chosen in $X$, and hence can be used for generating the root particles for $\mathfrak{A}_{U}$ and $\mathfrak{A}_{W}$, respectively. Now, suppose that for $k=0,1,2,\ldots$ we have $z_{k}\in\Lambda_{i}$ and $z^{\prime}_{k}\in\Omega_{i^{\prime}}$ such that
$i_{U^{*}}(z_{k})=i_{W^{*}}(z^{\prime}_{k})\;.$ (47)
For extending $\mathfrak{A}_{U}$ and $\mathfrak{A}_{W}$, we should generate $z_{k+1}$ and $z^{\prime}_{k+1}$ according to $\frac{\lambda_{z_{k},U}(\cdot)}{\deg_{U}(z_{k})}$ and $\frac{\lambda_{z^{\prime}_{k},W}(\cdot)}{\deg_{W}(z^{\prime}_{k})}$. By (47) and Lemma 6.3 we can couple the choice of $z_{k+1}$ and $z^{\prime}_{k+1}$ in a way that we have again (47).
To conclude, we have generated sequences $z_{0},z_{1},z_{2},\ldots$ and $z^{\prime}_{0},z^{\prime}_{1},z^{\prime}_{2},\ldots$ so that we have (47) for each $k$. In particular, that means that we have $(i_{U^{*}}(z_{k}))^{\ddagger}=(i_{W^{*}}(z^{\prime}_{k}))^{\ddagger}$. By the virtue of Lemma 6.4, we have $i_{(U^{*})^{\dagger}}(z_{k})=i_{(W^{*})^{\dagger}}(z^{\prime}_{k})$. Definition 4.2 tells us that $\dagger$ is invariant under multiplying a kernel by any nonzero scalar. Also, note that for an akernel $\Gamma$ and an element $x$ of its ground space, $i_{\Gamma}(x)$ depends only on the component of $\Gamma$ containing $x$. These facts combined give
$i_{U^{\dagger}}(z_{k})=i_{U\llbracket\Lambda_{i}\rrbracket^{\dagger}}(z_{k})=i_{W\llbracket\Omega_{i^{\prime}}\rrbracket^{\dagger}}(z^{\prime}_{k})=i_{W^{\dagger}}(z^{\prime}_{k})\;.$
Using Lemma 6.4 again, we get $(i_{U}(z_{k}))^{\ddagger}=(i_{W}(z^{\prime}_{k}))^{\ddagger}$, as was needed. ∎
Proof of Lemma 6.6.
The ancestral path in $\mathfrak{U}_{U}$ is generated in the same way as we generate the sequence $z_{0},z_{1},\ldots$ in the definition of $\mathfrak{A}_{U}$. That is, if we view each type $(\mathsf{anc},\cdot)$ through the lens of the operator $\big{(}i_{U}(\cdot)\big{)}^{\ddagger}$, we get the distribution of $\mathfrak{A}_{U}$. By Lemma 6.4, this is equivalent to viewing the types $(\mathsf{anc},\cdot)$ on the ancestral path through the lens of $i_{U^{\dagger}}(\cdot)$. Definition 1.5 tells us that we can construct $\mathfrak{U}_{U}$ from by attaching an independent branching process $\mathfrak{X}_{U^{\dagger}}(\cdot)$ to each vertex of type $(\mathsf{anc},\cdot)$ on its ancestral path. By Proposition 3.1(i), this is the same as attaching independent branching processes $\mathfrak{B}(i_{U^{\dagger}}(\cdot))=\mathfrak{B}((i_{U}(\cdot))^{\ddagger})$ (c.f. Lemma 6.4). That is, the distribution $\mathfrak{A}_{U}$ fully determines the distribution $\mathfrak{U}_{U}$.
We repeat this argument for $W$. We conclude that $\mathfrak{U}_{U}$ and $\mathfrak{U}_{W}$ have the same distribution. ∎
6.3 Proof of Lemma 6.2
We begin with a definition. Given an $L^{\infty}$-kernel $W$ on $(X,\mathcal{B},\mu)$ with positive minimum degree and a decomposition $X=\bigsqcup_{k\in I}X_{k}$ into connected components, let $W^{\heartsuit}$ be given by
$W^{\heartsuit}(x,y)=\begin{cases}\frac{W(x,y)\mu(X_{k})}{\int_{X_{k}\times X_{k}}W(x,y)\mathsf{d}\mu(x)\mathsf{d}\mu(y)}&\textrm{ if $x,y\in X_{k}$ for some$k\in I$,}\\ 0&\textrm{ otherwise.}\end{cases}$ (48)
It is easy to verify that $W^{\heartsuit}$ is an $L^{\infty}$-kernel with positive minimum degree.
Our proof of Lemma 6.2 has two main steps. First, we show that there is a transformation from $i_{W^{\dagger}}$ to $i_{W^{\heartsuit}}$ that depends only on the fractional isomorphism class of $W^{\dagger}$. This is the subject of Lemma 6.11, and a substantial part of this section is devoted to building tools for its proof. Lemma 6.11 allows us to show that $W^{\heartsuit}$ and $U^{\heartsuit}$ are fractionally isomorphic in the setting of Lemma 6.2. A quick application of Theorem 5.1 then allows us to conclude that $U$ and $W$ are piecewise projectively fractionally isomorphic.
Since the objects $i_{W^{\dagger}}(x)$ are elements of $\mathbb{P}^{c}$, it turns out to be useful to have for each $L^{\infty}$-akernel $W$ on $X$ a corresponding $L^{\infty}$-akernel $K[\nu_{W}]$ on $\mathbb{P}^{c}$. The following proposition combines generalizations of parts of Proposition 6.8 and Claim 6.9 in [11].
Proposition 6.7.
Suppose that $\nu$ is a probability measure on $\mathbb{P}^{c}$ such that $\nu=\nu_{W}$ for some $L^{\infty}$-akernel $W$. Then $\mu_{\alpha}$ is absolutely continuous with respect to $\nu$ for $\nu$-almost every $\alpha\in\mathbb{P}^{c}$, with the Radon–Nikodym derivative satisfying $0\leq\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}\leq\|W\|_{\infty}$, and there is an akernel $K[\nu]\in L^{\infty}(\mathbb{P}^{c}\times\mathbb{P}^{c},\nu\times\nu)$ such that $K[\nu](\alpha,-)=\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}$ for $\nu$-almost every $\alpha\in\mathbb{P}^{c}$.
Proof.
Fix $W$ such that $\nu=\nu_{W}$. Since $W$ is an $L^{\infty}$-akernel, for $\mu$-almost every $x\in X$ we have $W(x,y)\in[0,\|W\|_{\infty}]$ for $\mu$-almost every $y\in X$. To extract the desired Radon-Nikodym derivative $\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}$, we shall now recall [11, Corollary E.2].
Lemma 6.8 (Corollary E.2 in [11]).
Let $(X,\mathcal{B})$ and $(Y,\mathcal{D})$ be standard Borel spaces. Suppose that $\mu$ is a Borel probability measure and $f\colon X\to Y$ is a Borel function. Write $\mathcal{C}\in\Theta_{\mu}$ for the minimum relatively complete sub-sigma-algebra of $\mathcal{B}$ that makes $f$ measurable. Then for every $h\in L^{2}(X,\mathcal{C},\mu)$ there is a Borel map $g\colon Y\to\mathbb{R}_{+}$ such that $h(x)=(g\circ f)(x)$ for $\mu$-almost every $x\in X$.
Since by Lemma 4.9 $\mathcal{C}(W)$ is the minimum relatively complete sub-sigma-algebra of $\mathcal{B}$ that makes $i_{W}$ measurable, by Lemma 6.8 we obtain for $\mu$-almost every $x\in X$ a function $g_{x}\colon\mathbb{P}^{c}\to\left[0,\|W\|_{\infty}\right]$ such that
$\mathbf{E}[W(x,-)|\mathcal{C}(W)]=g_{x}\circ i_{W}$ (49)
holds $\mu$-almost everywhere. Take $\alpha\in\mathbb{P}^{c}$ such that $\alpha=i_{W}(x)$ for some $x\in X$ for which (49) holds. For $A\in\mathbb{P}^{c}_{n}$ we have
$\mu_{\alpha}(p_{n,\infty}^{-1}(A))=i_{W}(x)(n+1)(A)=\int_{i_{W,n}^{-1}(A)}W(x,y)\mathsf{d}\mu(y)$ $=\int_{i_{W,n}^{-1}(A)}\mathbf{E}[W(x,y)|\mathcal{C}(W)]\mathsf{d}\mu(y)$
$=\int_{i_{W}^{-1}(p_{n,\infty}^{-1}(A))}g_{x}\circ i_{W}(y)\mathsf{d}\mu(y)$
$=\int_{p_{n,\infty}^{-1}(A)}g_{x}(\beta)\mathsf{d}\nu(\beta)\;.$
Since $\mathcal{B}(\mathbb{P}^{c})$ is generated by $\bigcup_{n\in\mathbb{N}_{0}}\{p_{n,\infty}^{-1}(A):A\in\mathcal{B}(\mathbb{P}^{c}_{n})\}$, it follows that $g_{x}$ is the desired Radon-Nikodym derivative $\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}$.
We define a Borel probability measure $\tau$ on $\mathbb{P}^{c}\times\mathbb{P}^{c}$ as follows. Let $A\in\mathcal{B}(\mathbb{P}^{c}\times\mathbb{P}^{c})$ and write $A_{\alpha}=\{\beta\in\mathbb{P}^{c}:(\alpha,\beta)\in A\}$. Note that the assignment $\mathbb{P}^{c}\ni\alpha\mapsto\mu_{\alpha}(A_{\alpha})\in[0,c]$ is defined $\nu$-almost everywhere, and it is measurable as a consequence of [11, Claim 6.2]. Now set $\tau(A)=\int_{\alpha\in\mathbb{P}^{c}}\mu_{\alpha}(A_{\alpha})\mathsf{d}\nu(\alpha)$. Observe that
$\tau(A)=\int_{\alpha\in\mathbb{P}^{c}}\int_{\beta\in A_{\alpha}}\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}(\beta)\mathsf{d}\nu(\beta)\mathsf{d}\nu(\alpha)=\int_{(\alpha,\beta)\in A}\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}(\beta)\mathsf{d}\nu^{2}(\alpha,\beta)\;,$
so $\tau$ is absolutely continuous with respect to $(\nu\times\nu)$ with its Radon-Nikodym derivative $K[\nu]$ satisfying $K[\nu](\alpha,-)=\frac{\mathsf{d}\mu_{\alpha}}{\mathsf{d}\nu}$ for $\nu$-almost every $\alpha\in\mathbb{P}^{c}$ up to a $\nu$-nullset. ∎
The following lemma is a straightforward generalization of [11, Theorem 6.10] which formalizes the link between $W$ and $K[\nu_{W}]$. Given an $L^{\infty}$-akernel $W$ and a relatively complete sub-sigma-algebra $\mathcal{C}\in\Theta_{\mu}$, write $W_{\mathcal{C}}$ for $\mathbf{E}[W|\mathcal{B}\times\mathcal{C}]$.
Lemma 6.9.
For every $L^{\infty}$-akernel $W$ we have
$W_{\mathcal{C}(W)}(x,y)=K[\nu_{W}](i_{W}(x),i_{W}(y))$ (50)
for $(\mu\times\mu)$-almost every $(x,y)\in X\times X$.
The function $\deg_{W^{\heartsuit}}$ plays a key role since it relates $W^{\dagger}$ and $W^{\heartsuit}$. More specifically, we have
$W^{\dagger}(x,y)\deg_{W^{\heartsuit}}(y)=W^{\heartsuit}(x,y)$ (51)
for each $x,y\in X$. For an $L^{\infty}$-kernel $W$ on $(X,\mathcal{B},\mu)$ let $\chi_{W}$ be the measure on $(X,\mathcal{B})$ given by
$\mathsf{d}\chi_{W}=\deg_{W^{\heartsuit}}\mathsf{d}\mu$ (52)
and $\pi_{W}=(i_{W^{\dagger}})_{*}\chi_{W}$ be the pushforward measure on $\mathbb{P}^{c}$ ($c=\|W\|_{\infty}$) of $\chi_{W}$ via $i_{W^{\dagger}}$. Note that $\chi_{W}$ and subsequently also $\pi_{W}$ are probability measures. The following lemma, which is the key intermediate step in our proof of Lemma 6.11, says that the measures $\pi_{W}$ and $\pi_{U}$ are equal whenever the Markov renormalizations $W^{\dagger}$ and $U^{\dagger}$ are fractionally isomorphic. This is the one place where we apply the theory of Markov chains.
Lemma 6.10.
Suppose that $W$ and $U$ are $L^{\infty}$-kernels with positive minimum degree such that $W^{\dagger}$ and $U^{\dagger}$ are fractionally isomorphic. Write $c$ for $\|W^{\dagger}\|_{\infty}$. Then the measures $\pi_{W}$ and $\pi_{U}$ on $\mathbb{P}^{c}$ are equal.
Proof.
Write $\nu$ for $\nu_{W^{\dagger}}=\nu_{U^{\dagger}}$. Proposition 6.7 returns an akernel $K[\nu]$.
Before we dive into the proof, let us briefly motivate the steps in our proof. The basic idea is that we would like run a Markov chain on $\mathbb{P}^{c}$ with the transpose of $K[\nu]$ as its transition probability kernel, apply Theorem 4.1 to show that the Markov chain converges, and demonstrate that the unique limit probability measure is equal to both $\pi_{W}$ and $\pi_{U}$. The first significant hurdle we encounter is reducibility, that is, the existence of multiple connected components. We deal with this by splitting the space $\mathbb{P}^{c}$ into the connected components of $K[\nu]$, running a Markov chain in each of them, and then suitably amalgamating the Markov chains. The second obstacle is unsynchronized periodicity amongst countably infinitely many connected components. To handle this, we first analyze the behaviour for any finite number $M$ of connected components and then utilize the fact that the contribution of the remaining infinite tail necessarily vanishes as $M$ tends to infinity.
We begin by laying the groundwork for connected components in $K[\nu]$. Let $\mathcal{F}$ be the collection of Borel sets $B\subset\mathbb{P}^{c}$ such that
$\int_{B}\int_{\mathbb{P}^{c}\setminus B}K[\nu](\alpha,\beta)\mathsf{d}\nu(\alpha)\nu(\beta)=0\;.$ (53)
This formula strongly resembles the formula in Definition 2.1(iii) defining separation between connected components, and indeed the intended meaning is the same. But we must be more careful as $K[\nu]$ is an akernel; indeed, the notion of connected components is more problematic in the asymmetric case, just as it is in oriented graphs. Here the link between $K[\nu]$ and $W$ via $W^{\dagger}$ lets us draw upon structural information coming from connected components in $W$. In particular, we show in (A2) below that (53) is equivalent to $\int_{\mathbb{P}^{c}\setminus B}\int_{B}K[\nu](\alpha,\beta)\mathsf{d}\nu(\alpha)\nu(\beta)=0$. This will allow us to introduce the notion of ‘connected components of $K[\nu]$’.
Claim.
The following hold for all $B\in\mathcal{F}$.
  1. (A1)
    The set $i_{W^{\dagger}}^{-1}(B)$ is a (possibly empty) union of connected components of $W$, modulo a $\mu$-null set.
  2. (A2)
    We have $\mathbb{P}^{c}\setminus B\in\mathcal{F}$.
Proof.
Write $S$ for $i_{W^{\dagger}}^{-1}(B)$. Since $i_{W^{\dagger}}$ is $\mathcal{C}(W^{\dagger})$-measurable by Lemma 4.9, by (50) and (53) we have
$\int_{S}\int_{X\setminus S}W^{\dagger}(x,y)\mathsf{d}\mu(x)\mu(y)=\int_{S}\int_{X\setminus S}W^{\dagger}_{\mathcal{C}(W^{\dagger})}(x,y)\mathsf{d}\mu(x)\mu(y)=\int_{B}\int_{\mathbb{P}^{c}\setminus B}K[\nu](\alpha,\beta)\mathsf{d}\nu(\alpha)\nu(\beta)=0\;.$ (54)
Together with the definition of $W^{\dagger}$, this means that $W$ is zero $(\mu\times\mu)$-almost everywhere on $S\times(X\setminus S)$. Hence, (A1) follows. Now $W$ is symmetric, so both $W$ and $W^{\dagger}$ are zero $(\mu\times\mu)$-almost everywhere on $(X\setminus S)\times S$. By applying (50) we obtain
$\int_{\mathbb{P}^{c}\setminus B}\int_{B}K[\nu](\alpha,\beta)\mathsf{d}\nu(\alpha)\nu(\beta)=\int_{X\setminus S}\int_{S}W^{\dagger}_{\mathcal{C}(W^{\dagger})}(x,y)\mathsf{d}\mu(x)\mu(y)=\int_{X\setminus S}\int_{S}W^{\dagger}(x,y)\mathsf{d}\mu(x)\mu(y)=0\;,$
so (A2) follows. ∎
We say that two connected components of $W$ are inseparable if for every $B\in\mathcal{F}$ we have that $i_{W^{\dagger}}^{-1}(B)$, modulo a $\mu$-null set, contains both connected components or neither of them. Inseparability induces an equivalence relation on the collection of connected components of $W$; let $I$ be the index set for the countably many equivalence classes. Without loss of generality, we have $I=\mathbb{N}$ or $I=[N]$ for some $N\in\mathbb{N}$.
Let $k\in I$. For each $j\in I\setminus\{k\}$ pick $A_{k,j}\in\mathcal{F}$ such that $i_{W^{\dagger}}^{-1}(A_{k,j})$, modulo a $\mu$-null set, contains all of the connected components in the $k$th equivalence class and none of those in the $j$th equivalence class; this is possible by the definition of inseparability and the properties (A1) and (A2). Let $S_{k}:=\bigcap_{j\in I\setminus\{k\}}A_{k,j}$ and $X_{k}:=i_{W^{\dagger}}^{-1}(S_{k})$. We have $\nu(S_{k})=\mu(X_{k})$, and it is immediate from (53) that $S_{k}\in\mathcal{F}$. Note that $S_{k}$ serves as the $k$th connected component of $K[\nu]$ (c.f. discussion after (53)) and $X_{k}$ is the union of all the connected components in the $k$th equivalence class, modulo a $\mu$-null set.
We now prepare to run a Markov chain on $S_{k}$. Write $\nu_{k}$ and $\mu_{k}$ for the probability measures $\frac{\nu_{\restriction S_{k}}}{\nu(S_{k})}$ and $\frac{\mu_{\restriction X_{k}}}{\mu(X_{k})}$ respectively. Now write $W_{k}$ for $W_{\restriction X_{k}}$. Here $W_{k}$ is an $L^{\infty}$-kernel on $(X_{k}\times X_{k},\mu_{k}\times\mu_{k})$ with positive minimum degree, so by Lemma 4.3 $W_{k}^{\dagger}$ is an $L^{\infty}$-akernel on $(X_{k}\times X_{k},\mu_{k}\times\mu_{k})$ with positive minimum degree. Hence, $K[\nu]_{k}(\alpha,\beta):=\nu(S_{k})K[\nu]_{\restriction S_{k}}(\alpha,\beta)$ is an $L^{\infty}$-akernel on $(S_{k}\times S_{k},\nu_{k}\times\nu_{k})$ with positive minimum degree, say $\delta>0$, by Lemma 6.9 and Lemma 4.9. By modifying $K[\nu]_{k}$ on a $(\nu_{k}\times\nu_{k})$-null set, we may assume without loss of generality that $K[\nu]_{k}$ satisfies $\deg_{K[\nu]_{k}}(\alpha)\geq\delta$ for all $\alpha\in S_{k}$. Let $\Phi_{k}$ be a Markov chain on $S_{k}$ with transition probability kernel $Q_{k}(\alpha,A)=\int_{A}K[\nu]_{k}(\beta,\alpha)\mathsf{d}\nu_{k}(\beta)$ for $A\in\mathcal{B}(S_{k})$. We prepare to apply Theorem 4.1 to $\Phi_{k}$.
Claim.
The Markov chain $\Phi_{k}$ is $\nu_{k}$-irreducible.
Proof.
For $A\in\mathcal{B}(S_{k})$ and $\alpha\in S_{k}$ write $q_{\alpha,A}$ for the probability that $\Phi_{k}$ started at $\alpha$ satisfies $\Phi_{k}(n)\in A$ for some $n\in\mathbb{N}$. For the sake of contradiction, assume that there exists $A_{0}\in\mathcal{B}(S_{k})$ with $\nu_{k}(A_{0})>0$ such that $q_{\alpha,A_{0}}=0$ for some $\alpha\in S_{k}$. Let $R_{A_{0}}$ be the set of all such $\alpha\in S_{k}$. Because $K[\nu]_{k}$ has positive minimum degree and bounded $L^{\infty}$-norm, we have $\nu_{k}(R_{A_{0}}),\nu_{k}(S_{k}\setminus R_{A_{0}})>0$. Writing $Q_{k,\alpha}$ for the measure on $\mathcal{B}(S_{k})$ given by $Q_{k,\alpha}(A)=Q_{k}(\alpha,A)$, for $\alpha\in R_{A_{0}}$ we have
$0=q_{\alpha,A_{0}}=Q_{k,\alpha}(A_{0})+\int_{S_{k}\setminus(A_{0}\cup R_{A_{0}})}q_{\beta,A_{0}}\mathsf{d}Q_{k,\alpha}(\beta)\;.$ (55)
Since $q_{\beta,A_{0}}>0$ for $\beta\in S_{k}\setminus(A_{0}\cup R_{A_{0}})$, we get $Q_{k}(\alpha,S_{k}\setminus R_{A_{0}})=0$ from (55). Hence, we have $R_{A_{0}}\in\mathcal{F}$. But now $i_{W^{\dagger}}^{-1}(R_{A_{0}})$ and $i_{W^{\dagger}}^{-1}(S_{k}\setminus R_{A_{0}})$ are unions of connected components of $W$ with $R_{A_{0}},S_{k}\setminus R_{A_{0}}\in\mathcal{F}$ by (A1) and(A2), which contradicts the fact that $S_{k}\in\mathcal{F}$ represents an equivalence class of inseparable connected components in $W$. ∎
Observe that the pushforward measure $\pi_{k}$ of $\chi_{W_{k}}$ via $(i_{W^{\dagger}})_{\restriction X_{k}}$ is an invariant probability measure on $\mathcal{B}(S_{k})$. Indeed, since $i_{W^{\dagger}}$ is $\mathcal{C}(W^{\dagger})$-measurable by Lemma 4.9, by unpacking definitions we have
$\begin{split}\pi_{k}(A)=((i_{W^{\dagger}})_{\restriction X_{k}})_{*}\chi_{W_{k}}(A)&=\int_{i_{W^{\dagger}}^{-1}(A)}\deg_{W_{k}^{\heartsuit}}(x)\mathsf{d}\mu_{k}(x)\\ &=\int_{i_{W^{\dagger}}^{-1}(A)}\int_{X_{k}}W_{k}^{\dagger}(x,y)\deg_{W_{k}^{\heartsuit}}(y)\mathsf{d}\mu_{k}(y)\mathsf{d}\mu_{k}(x)\\ &=\int_{X_{k}}\int_{i_{W^{\dagger}}^{-1}(A)}W_{k}^{\dagger}(x,y)\mathsf{d}\mu_{k}(x)\deg_{W_{k}^{\heartsuit}}(y)\mathsf{d}\mu_{k}(y)\\ &=\int_{X_{k}}\int_{i_{W^{\dagger}}^{-1}(A)}\nu(S_{k})W^{\dagger}_{\mathcal{C}(W^{\dagger})}(x,y)\mathsf{d}\mu_{k}(x)\mathsf{d}\chi_{W_{k}}(y)\\ &=\int_{S_{k}}\int_{A}K[\nu]_{k}(\beta,\alpha)\mathsf{d}\nu_{k}(\beta)\mathsf{d}\pi_{k}(\alpha)=\int_{S_{k}}Q_{k}(\alpha,A)\mathsf{d}\pi_{k}(\alpha)\end{split}$ (56)
for every $A\in\mathcal{B}(S_{k})$. Furthermore, we define a full irreducibility probability measure $\psi_{k}$ for $\Phi_{k}$ by
$\psi_{k}(A):=\int_{S_{k}}\sum_{j=1}^{\infty}2^{-j-1}Q_{k}^{j}(\alpha,A)\mathsf{d}\nu_{k}(\alpha),\quad A\in\mathcal{B}(S_{k})\;.$
Hence, $\Phi_{k}$ is a positive recurrent Markov chain. Let $p_{k}$ be the period of $\Phi_{k}$. For each $n\in\mathbb{N}_{0}$ we define a measure $\pi_{k,n}$ on $\mathcal{B}(S_{k})$ given by $\pi_{k,n}(A)=\int_{S_{k}}\frac{1}{p_{k}}\sum_{r=0}^{p_{k}-1}Q_{k}^{np_{k}+r}(\alpha,A)\mathsf{d}\nu_{k}(\alpha)$. Since $W_{k}$ has positive minimum degree, the measures $\chi_{W_{k}}$ and $\mu_{k}$ are equivalent. Thus, the measures $\pi_{k}$ and $\nu_{k}$ are equivalent. By applying Theorem 4.1, we obtain
$\lim_{n\to\infty}\|\pi_{k,n}-\pi_{k}\|_{TV}=0\;.$ (57)
We now extend (57) from one component to multiple components. For periodicity reasons, we start with finitely many components. Let $M\in I$. Write $S_{(M)}$ and $X_{(M)}$ for $\bigcup_{k\in[M]}S_{k}$ and $\bigcup_{k\in[M]}X_{k}$ respectively. Let $\nu_{(M)}$ and $\mu_{(M)}$ denote the probability measures $\frac{\nu_{\restriction S_{(M)}}}{\nu(S_{(M)})}$ and $\frac{\mu_{\restriction X_{(M)}}}{\mu(X_{(M)})}$ respectively. Let $W_{(M)}(x,y):=W_{\restriction X_{(M)}}(x,y)$ and $K[\nu]_{(M)}(\alpha,\beta):=\nu(S_{(M)})K[\nu]_{\restriction S_{(M)}}(\alpha,\beta)$ be akernels on $(X_{(M)}\times X_{(M)},\mu_{(M)}\times\mu_{(M)})$ and $(S_{(M)}\times S_{(M)},\nu_{(M)}\times\nu_{(M)})$ respectively. Let $p_{(M)}$ be the least common multiple of $\{p_{i}:i\in[M]\}$. Let $Q_{(M)}(\alpha,A)=\int_{A}K[\nu]_{(M)}(\beta,\alpha)\mathsf{d}\nu_{(M)}(\beta)$ define a transition probability kernel. Let $\pi_{(M)}$ be the pushforward measure of $\chi_{W_{(M)}}$ via $(i_{W^{\dagger}})_{\restriction X_{(M)}}$. For each $n\in\mathbb{N}_{0}$ let $\pi_{(M),n}$ be a measure on $\mathcal{B}(S_{(M)})$ given by $\pi_{(M),n}(A)=\int_{S_{(M)}}\frac{1}{p_{(M)}}\sum_{r=0}^{p_{(M)}-1}Q_{(M)}^{np_{(M)}+r}(\alpha,A)\mathsf{d}\nu_{(M)}(\alpha)$.
By unpacking the definitions of $\chi_{W_{(M)}}$ and $\chi_{W_{k}}$, we get the identity
$\pi_{(M)}(A)=\sum_{k\in[M]}\tfrac{\nu(S_{k})}{\nu(S_{(M)})}\cdot\pi_{k}(A\cap S_{k})\;.$ (58)
By unpacking the definitions of $Q_{(M)}$, $Q_{k}$, $\nu_{(M)}$ and $\nu_{k}$, recalling the definition of the $n$-step transition probability kernel and applying the defining property of $\mathbb{P}^{c}\setminus S_{k}\in\mathcal{F}$ given in (53), we obtain for $\alpha\in S_{k}$ the identity $Q_{(M)}^{n}(\alpha,A)=Q_{k}^{n}(\alpha,A\cap S_{k})$. Hence, we have
$\pi_{(M),n}(A)=\sum_{k\in[M]}\tfrac{\nu(S_{k})}{\nu(S_{(M)})}\cdot\tfrac{p_{k}}{p_{(M)}}\sum_{\ell=1}^{p_{(M)}/p_{k}}\pi_{k,n+\ell-1}(A\cap S_{k})\;.$ (59)
Now by applying (58), (59) and the definition of the total variation norm, we get
$\|\pi_{(M),n}-\pi_{(M)}\|_{TV}\leq\sum_{k\in[M]}\tfrac{\nu(S_{k})}{\nu(S_{(M)})}\cdot\tfrac{p_{k}}{p_{(M)}}\sum_{\ell=1}^{p_{(M)}/p_{k}}\|\pi_{k,n+\ell-1}-\pi_{k}\|_{TV}\;.$
Hence, by (57) we have
$\lim_{n\to\infty}\|\pi_{(M),n}-\pi_{(M)}\|_{TV}=0\;.$ (60)
The argument above deals with the case when $I$ is finite, so it remains to consider the case $I=\mathbb{N}$. In this case, we run a Markov chain on the entire space. Without loss of generality, we may assume that $k\mapsto\mu(X_{k})$ is non-increasing. Let $Q(\alpha,A)=\int_{A}K[\nu](\beta,\alpha)\mathsf{d}\nu(\beta)$ define a transition probability kernel. The point is that our previous analysis for finitely many components is enough to control most of the behaviour, and we can easily control the residual tail of small measure. To make this precise, for $\varepsilon>0$ let $M_{\varepsilon}$ be the least $m\in\mathbb{N}$ such that $\nu(\mathbb{P}^{c}\setminus S_{(m)})\leq\varepsilon$ and $N_{\varepsilon}\in\mathbb{N}$ be such that for all $n\geq N_{\varepsilon}$ we have $\|\pi_{(M_{\varepsilon}),n}-\pi_{(M_{\varepsilon})}\|_{TV}\leq 2\varepsilon$; these are well defined because the sequence $(\nu(S_{k}))k\in\mathbb{N}$ is monotone decreasing with limit 00 and we have (60). For each $n\in\mathbb{N}_{0}$ and $M\in I$ let $\rho_{M,n}$ be a measure on $\mathcal{B}(\mathbb{P}^{c})$ given by $\rho_{M,n}(A)=\int_{\mathbb{P}^{c}}\frac{1}{p_{(M)}}\sum_{r=0}^{p_{(M)}-1}Q^{np_{(M)}+r}(\alpha,A)\mathsf{d}\nu(\alpha)$. For each $n\in\mathbb{N}_{0}$ let $\pi_{n}=\rho_{M_{2^{-n-2}},N_{2^{-n-2}}}$. By unpacking the definitions, we obtain
$\|\pi_{n}-\pi_{W}\|_{TV}\leq\left\|\pi_{(M_{2^{-n-2}}),N_{2^{-n-2}}}-\pi_{(M_{2^{-n-2}})}\right\|_{TV}+2\mu\left(\mathbb{P}^{c}\setminus S_{(M_{2^{-n-2}})}\right)\leq 2^{-n},$
so $\lim_{n\to\infty}\|\pi_{n}-\pi_{W}\|_{TV}=0$.
A similar line of reasoning for $U$ instead of $W$ yields $\lim_{n\to\infty}\|\pi_{n}-\pi_{U}\|_{TV}=0$, so we conclude that $\pi_{W}$ and $\pi_{U}$ are equal. ∎
The following lemma shows that the transformation from $i_{W^{\dagger}}$ to $i_{W^{\heartsuit}}$ depends only on the fractional isomorphism class of the Markov renormalization $W^{{\dagger}}$.
Lemma 6.11.
Suppose that $W$ and $U$ are $L^{\infty}$-kernels with positive minimum degree such that $W^{{\dagger}}$ and $U^{{\dagger}}$ are fractionally isomorphic. Then there is a function $f\colon\mathbb{P}^{c}\to\mathbb{P}^{c}$ such that $i_{W^{\heartsuit}}=f\circ i_{W^{\dagger}}$ and $i_{U^{\heartsuit}}=f\circ i_{U^{\dagger}}$.
Proof.
Write $\nu$ for $\nu_{W^{\dagger}}=\nu_{U^{\dagger}}$. Proposition 6.7 returns an akernel $K[\nu]$. Write $\pi$ for $\pi_{W}=\pi_{U}$, which is well-defined by Lemma 6.10. We shall inductively construct a sequence $(f_{k})_{k\in\mathbb{N}_{0}}$ of functions with the following properties. First, for all $k\in\mathbb{N}_{0}$ we have $f_{k}\colon\mathbb{P}^{c}\to\mathbb{P}^{c}_{k}$. Second, for $k\in\mathbb{N}$ and $j\in[k-1]_{0}$ we have $f_{k}(\alpha)(j)=f_{k-1}(\alpha)(j)$. Third, for all $k\in\mathbb{N}_{0}$ we have $i_{W^{\heartsuit},k}=f_{k}\circ i_{W^{\dagger}}$ and $i_{U^{\heartsuit},k}=f_{k}\circ i_{U^{\dagger}}$. Observe that the definitions of the functions $i_{W^{\heartsuit}}$ and $i_{U^{\heartsuit}}$ imply that the function $f\colon\mathbb{P}^{c}\to\mathbb{P}^{c}$ given by $f(\alpha)(k)=f_{k}(\alpha)(k)$ for $k\in\mathbb{N}_{0}$ has the desired properties.
For $k=0$ let $f_{0}\colon\mathbb{P}^{c}\to\mathbb{P}^{c}_{0}$ be given by $f_{0}(\alpha)=\star$. This function has the desired properties because $i_{W^{\heartsuit},0}$ and $i_{U^{\heartsuit},0}$ are both constant functions mapping to $\star$. For $k\in\mathbb{N}$ we set $f_{k}(\alpha)(j)=f_{k-1}(\alpha)(j)$ for $j\in[k-1]_{0}$ and define $f_{k}(\alpha)(k)$ as follows. For measurable $A\subset\mathbb{P}^{c}_{k-1}$ set
$f_{k}(\alpha)(k)(A)=\int_{f_{k-1}^{-1}(A)}K[\nu](\alpha,\beta)\mathsf{d}\pi(\beta)\;.$
We shall now verify that $f_{k}$ has the requisite properties. The first two are immediate, so we shall focus on the third one. Take $x\in X$. The domain of the functions $i_{W^{\heartsuit},k}$, $i_{U^{\heartsuit},k}$ and $f_{k}\circ i_{U^{\dagger}}$ has components indexed by $[k]_{0}$. We first verify the claimed identity for components indexed by $j\in[k-1]_{0}$. This claim easily follows by induction. Indeed, we have
$(f_{k}\circ i_{W^{\dagger}})(x)(j)=(f_{k-1}\circ i_{W^{\dagger}})(x)(j)=i_{W^{\heartsuit},k-1}(x)(j)=i_{W^{\heartsuit},k}(x)(j)\;,$
as was needed. We now turn to the case $j=k$. For an arbitrary measurable $A\subset\mathbb{P}^{c}_{k-1}$ we have
$f_{k}(i_{W^{\dagger}}(x))(k)(A)$ $=\int_{f_{k-1}^{-1}(A)}K[\nu](i_{W^{\dagger}}(x),\beta)\mathsf{d}\pi(\beta)=\int_{i_{W^{\dagger}}^{-1}(f_{k-1}^{-1}(A))}W^{\dagger}_{\mathcal{C}(W^{\dagger})}(x,y)\mathsf{d}\chi_{W}(y)$
Lemma 4.9, $i_{W^{\heartsuit},k-1}=f_{k-1}\circ i_{W^{\dagger}}$ and (52) $=\int_{i_{W^{\heartsuit},k-1}^{-1}(A)}W^{\dagger}(x,y)\deg_{W^{\heartsuit}}(y)\mathsf{d}\mu(y)$
by (51) $=\int_{i_{W^{\heartsuit},k-1}^{-1}(A)}W^{\heartsuit}(x,y)\mathsf{d}\mu(y)=i_{W^{\heartsuit},k}(x)(k)(A)\;,$
as was needed. Hence, we have $i_{W^{\heartsuit},k}=f_{k}\circ i_{W^{\dagger}}$. An analogous line of reasoning yields $i_{U^{\heartsuit},k}=f_{k}\circ i_{U^{\dagger}}$, thereby completing the proof. ∎
We conclude the section with the proof of Lemma 6.2.
Proof of Lemma 6.2.
Let $W$ and $U$ be $L^{\infty}$-kernels with positive minimum degree such that $W^{\dagger}$ and $U^{\dagger}$ are fractionally isomorphic. Lemma 6.11 returns a function $f\colon\mathbb{P}^{c}\to\mathbb{P}^{c}$ such that $i_{W^{\heartsuit}}=f\circ i_{W^{\dagger}}$ and $i_{U^{\heartsuit}}=f\circ i_{U^{\dagger}}$. For an arbitrary measurable $A\subset\mathbb{P}^{c}$ we have
$\nu_{W^{\heartsuit}}(A)=\mu(i_{W^{\heartsuit}}^{-1}(A))=\mu(i_{W^{\dagger}}^{-1}(f^{-1}(A)))=\nu_{W^{\dagger}}(f^{-1}(A))$
and analogously we have $\nu_{U^{\heartsuit}}(A)=\nu_{U^{\dagger}}(f^{-1}(A))$. Since $W^{{\dagger}}$ and $U^{{\dagger}}$ are fractionally isomorphic, we have $\nu_{W^{\dagger}}=\nu_{U^{\dagger}}$ and so $\nu_{W^{\heartsuit}}(A)=\nu_{U^{\heartsuit}}(A)$. Hence, $W^{\heartsuit}$ and $U^{\heartsuit}$ are fractionally isomorphic.
Now fix a decomposition $X=\bigsqcup_{i\in I}X_{i}$ into the countably many connected components of $W$. Fix an analogous decomposition $X=\bigsqcup_{j\in J}Y_{j}$ into the countably many connected components of $U$. In particular, by Theorem 5.1, we have for every connected bounded degree kernel $\Gamma^{*}$ that
$\sum_{i\in J:\text{$U^{\heartsuit}\llbracket Y_{i}\rrbracket$ is frac.\ iso.\ to $\Gamma^{*}$}}\mu(Y_{i})=\sum_{i\in I:\text{$W^{\heartsuit}\llbracket X_{i}\rrbracket$ is frac.\ iso.\ to $\Gamma^{*}$}}\mu(X_{i})\;.$
[…]
$\sum_{i\in J:\text{$\mu(Y_{i})(U^{\heartsuit})_{\restriction Y_{i}}$ is frac.\ iso.\ to $\Gamma^{*}$}}\mu(Y_{i})=\sum_{i\in I:\text{$\mu(X_{i})(W^{\heartsuit})_{\restriction X_{i}}$ is frac.\ iso.\ to $\Gamma^{*}$}}\mu(X_{i})\;.$ (61)
Claim.
Let $\Gamma$ be an arbitrary connected bounded degree kernel. Then for each $i\in I$, the following are equivalent.
  1. (i)
    $W_{\restriction X_{i}}$ is projectively fractionally isomorphic to $\Gamma$,
  2. (ii)
    $(W^{\heartsuit})_{\restriction X_{i}}$ is projectively fractionally isomorphic to $\Gamma$,
  3. (iii)
    $\mu(X_{i})\|\Gamma\|_{1}\cdot(W^{\heartsuit})_{\restriction X_{i}}$ is fractionally isomorphic to $\Gamma$, and
Proof.
The equivalence between (i) and (ii) is obvious, since $W_{\restriction X_{i}}$ and $(W^{\heartsuit})_{\restriction X_{i}}$ differ just by a scalar multiple. The direction (iii)$\Rightarrow$(ii) is obvious. For the direction (ii)$\Rightarrow$(iii), assume that $(W^{\heartsuit})_{\restriction X_{i}}$ is projectively fractionally isomorphic to $\Gamma$. That is $c\cdot(W^{\heartsuit})_{\restriction X_{i}}$ is fractionally isomorphic to $\Gamma$, for some $c>0$. It is our task to show that $c=\mu(X_{i})\|\Gamma\|_{1}$. Since two fractionally isomorphic kernels must have the same $L^{1}$-norm, we have $c\cdot\|(W^{\heartsuit})_{\restriction X_{i}}\|_{1}=\|\Gamma\|_{1}$. Using (48) (and also realizing that the restriction operation $\restriction X_{i}$ reweights the square $X_{i}\times X_{i}$ by a factor $\mu(X_{i})^{-2}$), we have $\|(W^{\heartsuit})_{\restriction X_{i}}\|_{1}=\frac{1}{\mu(X_{i})}$. ∎
To finish the proof, we need to show that for every connected bounded-degree kernel $\Gamma$, we have
$\sum_{i:\text{$U_{\restriction Y_{i}}$ is proj.frac.\ iso.\ to $\Gamma$}}\mu(Y_{i})=\sum_{i:\text{$W_{\restriction X_{i}}$ is proj.frac.\ iso.\ to $\Gamma$}}\mu(X_{i})\;.$
By the above Claim (and the obvious counterpart of this claim for the kernel $U$), this is equivalent to
$\sum_{i:\text{$\mu(Y_{i})\|\Gamma\|_{1}\cdot(U^{\heartsuit})_{\restriction Y_{i}}$ is frac.\ iso.\ to $\Gamma$}}\mu(X_{i})=\sum_{i:\text{$\mu(X_{i})\|\Gamma\|_{1}\cdot(W^{\heartsuit})_{\restriction X_{i}}$ is frac.\ iso.\ to $\Gamma$}}\mu(X_{i})\;.$
This follows from (61) for $\Gamma^{*}:=\frac{\Gamma}{\|\Gamma\|_{1}}$. ∎
7 Concluding remarks
Our first main result, Theorem 1.3, provides a characterization of akernels that yield the same distribution of the corresponding Bollobás–Janson–Riordan branching processes, under the assumption of bounded degrees. We do not know how much can this assumption be relaxed. More specifically, the definition of the branching process $\mathfrak{X}_{W}$ is sensible for every kernel $W$ for which $\deg_{W}(x)<\infty$ for almost every $x\in X$. The theory developed in [7, 6] applies to all $L^{1}$-kernels. [^6] A curious question is which results from  [7, 6] extend to kernels whose almost all degrees are finite. However, our Theorem 1.3 is proven only for akernels of bounded degree, a class somewhat broader than akernels of bounded $L^{\infty}$-norm (c.f. Footnote 4). As we explain in Remark 2.2, the reason why we were not able to extend Theorem 1.3 to $L^{1}$-kernels is our reliance on the Stone–Weierstraß Theorem. We leave the extension to $L^{1}$-kernels, or a counterexample in this setting as an open problem. Note that the ()($\Leftarrow$)( ⇐ )-direction, as we formulated it in Proposition 3.1, extends easily.
Our second main result, Theorem 6.1, is proven for $L^{\infty}$-kernels of positive minimum degree, whereas maximum potential generality seems to be nondegenerate $L^{1}$-kernels. There are three places where we rely on these degree assumptions. The first one is that we need the corresponding Markov renormalization be of bounded degree (per Lemma 4.3) so that we are in the setting of Theorem 1.3. We discussed a possible extension of Theorem 1.3 above. The second one is the extinction Lemma 4.7. It could be that Lemma 4.7 can be extended to the needed setting quite easily. Again, the proof of the ()($\Leftarrow$)( ⇐ ) direction, as we formulated it in Section 6.2, extends easily. The third place concerns invariant sub-sigma-algebras, introduced in [11] and recalled in Section 4.4. This theory was worked out in [11] by Grebík and Rocha in the context of $L^{2}$-kernels. In a personal communication, Grebík suggested that the correct generality may be that of $L^{1}$-kernels.
Appendix A Summary of notation
$(X,\mathcal{B},\mu)$ ground space for all kernels and akernels, see p.1.1
$\deg_{W}(x)$ degree of a akernel at $x\in X$ (measured on the first coordinate), see p.1.1
$\operatorname{deg^{min}}(W)$, $\operatorname{deg^{max}}(W)$ minimum and maximum degree of $W$, see p.1.1
$\mathfrak{X}_{W}$ Bollobás–Janson–Riordan branching process from akernel $W$, see p.1.1
$\mathfrak{U}_{W}$ uniform spanning tree branching process from kernel $W$, see p.1.5
$\mathfrak{U}_{W}^{-}$ descendant part of $\mathfrak{U}_{W}$, see p.1
$\mathbb{P}_{n}^{c}$, $\mathbb{P}^{c}$ spaces of iterated degree measures, see p.3, eq.(3) and p.6, eq.(6)
$p_{\ell,k}$ projection from $\mathbb{P}_{k}$ to $\mathbb{P}_{\ell}$, see p.4
$\mathcal{B}(\mathbb{P}^{c})$, $\mathcal{B}(\mathbb{P}_{n}^{c})$ Borel sets on $\mathbb{P}^{c}$ and $\mathbb{P}^{c}_{n}$, see p.2.2
$D(\alpha)$ ‘total mass’ of $\alpha\in\mathbb{P}^{c}$, see p.2.2
$\mu_{\alpha}$ lifting of an element $\alpha\in\mathbb{P}^{c}$ to an element in $\mathcal{M}_{\leq c}(\mathbb{P}^{c})$, see p.2.2
$i_{W,n}(x)$, $i_{W}(x)$ degree information at vertex $x$, see p.2.2
$\nu_{W}$ iterated degree measure, see p.2.2
$\mathfrak{B}(\alpha)$ branching process from iterated degree measure $\alpha\in\mathbb{P}^{c}$, see p.3.1
$\mathfrak{B}_{W}$ branching process $\mathfrak{B}(\alpha)$, where $\alpha$ is sampled from akernel $W$ using $i_{W}$, see p.3.1
$T^{\uparrow}$ introducing a new root preceding the old root, see p.3.2
$T_{1}\oplus T_{2}$ pasting two rooted trees along the root, see p.3.2
$\mathcal{T}_{n}$ trees of height at most $n$, see p.3.2
$\widehat{\nu}$ exponential tilting of measure $\nu$, see p.3.3
$W^{\dagger}$ Markov renormalization of a kernel, see p.4.2
$e_{F}$ coefficient associated to a tree, see p.14, eq.(14)
$f_{T,n}$, $\overline{f_{T,n}}$ tree functions, see p.3.2, Definition 3.2
$\mathcal{F}_{n}$, $\overline{\mathcal{F}_{n}}$ collections of functions $f_{T,n}$ and $\overline{f_{T,n}}$, see p.3.2
$\mathcal{C}^{W}_{n}$, $\mathcal{C}^{W}$ pullback sigma-algebras $\mathbb{P}^{c}_{n}$ via $i_{W,n}$ or of $\mathbb{P}^{c}$ via $i_{W}$, see p.4.4
$U_{\restriction Y}$ restriction of a kernel, see p.5
$U\llbracket Y\rrbracket$ restricted and rescaled kernel, see p.5
$\alpha^{\ddagger}$ degree tilting, see p.6.2
$\lambda_{W,x}$ neighborhood measure of $x$ in $W$, see p.6.2
[…]
$W^{\heartsuit}$ a renormalization inside components of $W$, see p.48
$K[\nu]$ the simplest akernel whose iterated degree measure is $\nu$, see p.6.7
$\chi_{W}$ probability measure on $X$, $\mathsf{d}\chi_{W}=\deg_{W^{\heartsuit}}\mathsf{d}\mu$, see p.51
$\pi_{W}$ probability measure on $\mathbb{P}^{C}$, $\pi_{W}=(i_{W^{\dagger}})_{*}\chi_{W}$, see p.52
$K[\nu]_{k}$ $=(K[\nu])\llbracket S_{k}\rrbracket$, see p.Claim
Contact details
Postal address: Institute of Computer Science of the Czech Academy of Sciences
Pod Vodárenskou věží 2
182 00, Prague
Czechia
Email: hladky@cs.cas.cz, hng@cs.cas.cz, limbach@cs.cas.cz